### v1

# LowDiff: Efficient Frequent Checkpointing via Low-Cost Differential for High-Performance Distributed Training Systems

Abstract
Distributed training of large deep-learning models often leads to failures, so checkpointing is commonly employed for recovery. State-of-the-art studies focus on frequent checkpointing for fast recovery from failures. However, it generates numerous checkpoints, incurring substantial costs and thus degrading training performance. Recently, differential checkpointing has been proposed to reduce costs, but it is limited to recommendation systems, so its application to general distributed training systems remains unexplored.
This paper proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost. Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently. It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance. We further enhance LowDiff with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression. Experiments on various workloads show that LowDiff can achieve checkpointing frequency up to per iteration with less than 3.1% runtime overhead.
I Introduction
Deep learning has demonstrated remarkable capabilities in the fields of computer vision [49, 20], natural language processing [14, 6] and so on. Recently, the advent of large models like ChatGPT [37] and GPT-4 [38] has captivated both academic and industrial attention. To efficiently train the large models, distributed training systems [27, 26, 46, 44] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations. Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [55]. For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [23]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [34].
To handle failures, checkpointing, which regularly writes the model state to storage (e.g., disks) for failure recovery, can confine the potential loss of the training progress to the interval between consecutive checkpoints, only requiring the recovery of the remaining progress. Traditional checkpointing saves the model states (including the model parameters and optimizer parameters) at epoch boundaries, but an epoch often runs for hours, so the epoch-level checkpointing often results in hours of GPU computation wasted on average for every failure recovery [36].
Consequently, state-of-the-art checkpointing studies on fast recovery (e.g., Checkfreq [36] and Gemini [54]) aim to increase the frequency of checkpointing from epochs to iterations (we call frequent checkpointing). Nevertheless, the frequent checkpointing generates a considerable number of checkpoints, whose creation and transmission will incur substantial computational and communication costs, and thus hinder the training process significantly. Recent studies on reducing such costs (e.g., Check-N-Run [15]) introduce differential checkpointing (DC for short), which only tracks and checkpoints the modified part of the model to reduce the checkpoint size, into recommendation systems by leveraging its unique sparse features.
However, when deploying DC with high frequency in general distributed training systems, we observe two challenges arising: 1) without the help of sparse features, frequent DC incurs significant computation cost required for compressing the original differentials into smaller ones (i.e., compressed differentials), thus hindering the training process, and 2) frequent DC causes significant training stalls, as frequent checkpoint writes incur substantial transmissions which can not be overlapped with training, thus degrading the training performance (see §III-A for details).
Fortunately, we also have two findings from the gradient compression techniques [7, 18, 35, 9, 50] widely used in distributed training systems for communication efficiency: 1) the compressed gradients can be reused to construct the compressed differentials for general deep-learning models. 2) Using compressed gradients eliminates the data dependency between training and checkpointing, enabling them to run in parallel. Moreover, the size of compressed gradients is significantly smaller than that of compressed differentials (see §III-B for details).
Therefore, to address the above challenges, the corresponding findings inspire us to leverage the compressed gradients in distributed training systems for low-cost differential checkpointing, so as to improve the overall training performance. Our motivation is that the compressed gradients can act as the compressed differentials with no DC computation cost and with a lower DC transmission cost.
In this paper, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials). Further, to enhance LowDiff, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
In scenarios without gradient compression, we further investigate how to efficiently reuse gradients for frequent DC construction and propose LowDiff+. In this case, directly compressing differentials would incur high computation costs, so we reuse and snapshot gradients as they are generated layer by layer during the backward pass, overlapping their offloading with computation, thus enabling a fine-grained reusing and snapshotting approach. Meanwhile, the model state is reconstructed in CPU memory by the reused gradients and regularly persisted to storage, both in parallel with ongoing GPU training, thus enabling high-frequency in-memory checkpointing and asynchronous persistence.
Our contributions include:
  • We observe two challenges in computation and transmission when deploying differential checkpointing directly in general distributed training systems with high frequency. Motivated by the two corresponding findings of compressed gradients, we propose LowDiff, which reuses compressed gradients to act as the compressed differential checkpoints, thereby lowering the cost (§III).
  • We design LowDiff, which realizes the reuse of compressed gradients via a queue structure and facilitates zero-copy transmission between the training and checkpointing processes. Atop LowDiff, we propose a batched gradient writing optimization by merging compressed gradients in CPU memory before persisting to storage. We also design a checkpointing configuration optimization to minimize the wasted GPU time (§IV).
  • We further propose LowDiff+ for efficient gradient reuse without compression. By exploiting the layer-by-layer generation of gradients during training, we design a layer-wise gradient reusing and snapshotting approach. We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§V).
  • We implement LowDiff atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff. Experiments on various models show that LowDiff can achieve checkpointing as frequently as once per iteration with less than 3.1% runtime overhead. Compared to the state-of-the-art methods, LowDiff can shorten training time by up to $89.2\%$ under high checkpointing frequencies (§VI & §VII).
II Background and Related Work
II-A Basics of Distributed Training Systems
A deep neural network (DNN) contains many model parameters. Training DNN models is the process of determining the optimal set of weights and biases that minimizes the loss function of the model, which measures the difference between the model predictions and the target values. The training process starts with a set of parameters and proceeds iteratively over a dataset many times (i.e., epochs), each of which contains multiple iterations, with an optimizer being used to adjust the parameters based on the gradients (i.e., derivatives of the loss function w.r.t. model parameters)
To accelerate the training process of large models and datasets [56, 28, 46, 44], distributed training systems are commonly adopted to distribute computations and data across multiple GPUs. Specifically, distributed training has four steps for the $t$-th iteration:
  • Forward pass. The model function is applied to the input data set to obtain the prediction.
  • Backward pass. A loss function is used to calculate the gradient on each worker.
  • Gradient synchronization. The gradients of different workers are synchronized through communication primitives.
  • Model update. The optimizer (e.g., Adam [24]) uses the synchronized gradient $G_{t}$ to update the model state (denoted by $M_{t}$), which contains the model parameters $x_{t}$ and the optimizer parameters $o_{t}$ (i.e., $M_{t}=(x_{t},o_{t})$), satisfying that
$M_{t+1}=M_{t}+Adam(G_{t}).$ (1)
Here, we specially introduce the widely used optimizer, Adam. The Adam optimizer maintains the first-order and second-order moments (i.e., the mean and variance of the gradients), and each has the same size as the model parameters, denoted by $\Psi$. Therefore, Adam requires an additional $2\Psi$ parameter storage beyond the model itself, which will support one of our insights in §III-B.
II-B Checkpointing Techniques
When training large models, failures frequently occur in distributed training systems due to extensive computational resources and long training times. To handle frequent training failures, checkpointing, a widely used technique in common training frameworks (e.g., PyTorch [39] and Tensorflow [1]), has been extensively studied [15, 36, 54, 19, 53, 8, 32]. It regularly copies the model state to local storage or remote storage during training. During recovery from the failure, all GPUs load the latest checkpoint and rerun the remaining process until the point of failure. To evaluate checkpointing performance, recent studies (e.g., [19]) define the wasted time metric as the sum of the recovery time from the latest checkpoint and the steady-state overhead (i.e., the GPU time for checkpointing when no failures have happened).
Frequent checkpointing. Traditional checkpointing is performed at the granularity of epochs, but an epoch may run for hours, thus incurring substantial wasted time for every failure [36]. Thus, state-of-the-art studies [36, 54] that aim for fast recovery increase the checkpointing frequency from epochs to iterations, which we call frequent checkpointing in this paper. The state-of-the-art systems are specified as follows:
Checkfreq [36] decouples checkpointing into snapshot and persist operations, and pipelines them with computation to enable frequent checkpointing. Specifically, the snapshot operation copies model parameters from GPU to CPU memory, while the persist operation asynchronously writes the snapshot to persistent storage. This design allows Checkfreq to create a checkpoint every 14–19 iterations.
Gemini [54] improves frequent checkpointing with the help of the CPU memory of the host machine and proposes a checkpointing traffic scheduling algorithm to mitigate the impact on model training. In this way, Gemini increases the checkpoint frequency over Checkfreq by up to $8\times$, even per-iteration frequency.
PCcheck [52] enables checkpointing as frequently as every 10 iterations by checkpointing to persistent main memory (PMEM). PCcheck utilizes concurrent checkpoints and minimizes time per checkpoint using pipelining techniques and multiple threads.
Just-in-time checkpointing [19] diverges from prior frequent checkpointing techniques. Instead of saving model states at regular intervals, Just-in-time checkpointing initiates a checkpoint only when a failure is detected and assumes redundancy across workers to ensure fault tolerance.
Despite the fast recovery of frequent checkpointing, the frequent creation and transmission of checkpoints will incur substantial costs, thus degrading the training performance. In this paper, our main goal is to lower the costs of frequent checkpointing for high-performance training.
Differential checkpointing. A recent study, Check-N-Run [15], on reducing the checkpointing costs introduces differential checkpointing (DC for short), which records the part of the model that is modified only during the last checkpoint interval so as to reduce the checkpoint size.
Specifically, DC saves its model state as full checkpoint ($C^{F}$) regularly. DC also saves its differential checkpoint of the $t$-th iteration (denoted by $C^{D}_{t}$) at an iteration-level high frequency. Then the model state $M_{n}$ at the $n$-th iteration can be reconstructed by merging the full and all differential checkpoints as
$M_{n}=C^{F}+C^{D}_{1}+C^{D}_{2}+\dots+C^{D}_{n}.$ (2)
However, we note that Check-N-Run implements DC only in recommendation systems, since the recommendation model has a unique property that its embedding table has a high sparsity, such that DC is particularly well suited for recommendation models where only a small part of the model parameters are updated after each iteration.
In contrast, general DNN models have to be updated entirely after each iteration since gradients are computed for all the model parameters, so in this case, deploying DC has to highly compress the differentials, leading to substantial computational cost, while higher frequency incurs more cost (see §III-A).
In this paper, we focus on how to enable low-cost DC for general DNN models, so as to extend the DC technique from recommendation systems to general distributed training systems.
II-C Gradient Compression Techniques
Gradient compression techniques have been widely used in high-performance distributed training systems [3, 16, 17, 30, 31, 47, 48, 51, 13, 4, 9, 50] to significantly reduce the communication overhead of the gradient synchronization (see §II-A). Two representative compression approaches are Sparsification [3, 51], which selects only a subset of the elements of the gradient, and Quantization [13, 5], which reduces the number of bits of each element of the gradient.
In this paper, we will observe that the costs of deploying DC in distributed training can be largely reduced by leveraging gradient compression (sparsification and quantization), which motivates our main idea (see §III-C for details).
III Observations and Motivation
(a) Computation Frequency
(b) Transmission Frequency
Figure 1: Impacts of DC computation and transmission frequency (in iterations) on training performance of GPT2-L.
In this section, we present two challenges of deploying frequent differential checkpointing (DC) in distributed training systems (§III-A) and discuss two corresponding findings from the gradient compression technique (§III-B). We also provide an analysis of the data dependency between training and checkpointing. Finally, we demonstrate how these observations motivate our main idea to address the challenges (§III-C).
III-A Challenges of DC in General Training
As stated in §II-B, DC may incur high costs when deployed in general distributed training systems without the help of the sparsity feature in recommendation systems, so we conduct measurements to observe that show the challenges of DC in general distributed training systems, especially under high frequency. We conduct our measurement on eight NVIDIA A100 GPUs with a 25Gbps network [54]. We evaluate a typical large-model training task GPT2-L [41], with a common differential checkpointing scheme (see Equation (2)). Detailed setups are shown in §VII-A.
Challenge 1: Frequent DC incurs significant computation cost for highly compressing differentials. In the absence of the sparsity feature’s support, DC has to apply the highly compressed scheme (compression ratio $\rho$ = 0.01 [15]) to the differentials (called compressed differentials), thus degrading the training performance. To measure the impact of compression computation cost on training, Figure 1(a) shows the training time of cases without compression and with different compression frequencies (e.g., 8 iteration compression frequency means the compression is performed every 8 iterations). We see that compared to the case without compression, the case with compression slows down the training process by $13\%\sim 57\%$, while higher frequency incurs slower training.
Challenge 2: Frequent DC causes significant transmission cost due to frequent checkpoint writes. Frequent DC needs to write the differential of model state to storage frequently, which blocks the training process frequently and incurs substantial transmission overhead, since the model update has to wait for the snapshotting of checkpoint [36, 53]. To measure the impact of transmission cost on training in terms of frequency, Figure 1(b) shows the training time of cases without differential transmission and with different differential transmission frequencies. We observe that compared to the case without differential transmission, the case with differential transmission slows down the training process by $12\%\sim 54\%$, while higher frequency incurs slower training.
Figure 2: Motivating example of LowDiff.
III-B Findings from Gradient Compression
To address the above challenges, we resort to the gradient compression technique, which is widely studied and deployed in both academia and industry (§II-C), and obtain the following two findings.
Finding 1: The compressed gradients can directly construct the compressed differentials for general DNN models. Gradients of a general DNN model are used to adjust the model state via the optimizer (e.g., Adam), so based on Equation (1), we have $Adam(G_{t})=M_{t+1}-M_{t}$. Meanwhile, the differential checkpoint of the $t$-th iteration $C_{t}^{D}$ also equals $M_{t+1}-M_{t}$ by definition (see §II-B). Thus, we have $C_{t}^{D}=Adam(G_{t})$. Therefore, the compressed version of differential checkpoint $C_{t}^{D}$ can be directly constructed via the optimizer by the compressed gradient $\widetilde{G}_{t}$, which can be obtained in advance by the gradient compression operation.
Finding 2: The compressed gradients are one-third of the compressed differentials in size. As mentioned in §II-A, if the model parameters are of size $\Psi$, the Adam optimizer parameters occupy $2\Psi$, making a full checkpoint $3\Psi$ in total. Next, Equation (2) implies that a differential checkpoint has the same size as the full one, i.e., $3\Psi$, which includes the model parameters and the Adam optimizer parameters. By contrast, the size of a gradient is equal to that of the model parameters, i.e., $\Psi$. Therefore, applying the same compression method results in a compressed gradient three times smaller than a compressed differential checkpoint.
III-C Main idea
Challenges 1 and 2 show that frequent DC has high computation and transmission cost, including compression stalls (caused by DC’s differential compression) and transmission stalls (caused by DC’s writes). Figure 2(a) illustrates that during a 3-iteration training, DC has three compression stalls and three transmission stalls, hindering the training process.
Findings 1 and 2 show that the compressed gradient can construct the compressed differentials and its size is much smaller than the compressed differential. Thus, the compressed gradient can be leveraged to act as the compressed differential checkpoint of DC, so as to 1) remove the compression stalls via eliminating DC’s differential compression, and 2) reduce or even remove the transmission stalls since the small-size compressed gradient is likely to enable DC’s transmission to be overlapped with the training process.
Therefore, we propose LowDiff, a low-cost differential checkpointing framework, whose main idea is to reuse the compressed gradients to act as the differential checkpoints with no DC compression cost and with a lower DC transmission cost. Figure 2(b) illustrates that during a 3-iteration training, LowDiff has no compression and transmission stalls by reusing $\widetilde{G}_{t}$, $\widetilde{G}_{t+1}$ and $\widetilde{G}_{t+2}$ as three differential checkpoints, speeding up the training.
III-D Data dependency analysis for LowDiff
Requirement of LowDiff. LowDiff requires that gradient reuse be decoupled from training, allowing DC to run in parallel. Without this decoupling, gradient reuse may introduce additional stalls in training (i.e., data dependency [53]). Specifically, in standard training, checkpointing can run independently of the forward and backward passes. However, checkpointing is data dependent on the model update pass [53], since it must read the parameter after model updates (i.e., Write-after-Read (WAR) [21]), which incurs a training stall for synchronization. Prior work [36, 15] also takes the data dependencies into account to design a checkpointing strategy that overlaps with training and reduces stalls. Consequently, it is crucial to carefully examine the coupling introduced by gradients between training and checkpointing. Next, we first show the data dependency when deploying DC directly, and then explore how reusing compressed gradients for DC interacts with the training process.
Deploying DC directly introduces additional data dependencies. As shown in Figure 3(a), checkpointing depends not only on the latest model state but also on the model state from $n$ iterations earlier. This introduces two major issues: 1) The model update in iteration $t+1$ must wait for checkpointing in iteration $t$, leading to a training stall due to the transmission of checkpointing. 2) The model state from iteration tnt-n must be retained in GPU memory until the differential computation in iteration $t$, resulting in extra GPU memory overhead. Moreover, both checkpoint compression and writing lie on the critical path of these data dependencies, and therefore cannot be effectively overlapped with training.
Reusing compressed gradients for DC eliminates such data dependency. Gradients are generated during the backward pass and remain unchanged thereafter. The update pass only reads these gradients (read-only) to update the model state, which removes the Write-after-Read (WAR) dependency and allows gradient reuse to be performed immediately after backpropagation without waiting for the model update. Moreover, by treating the compressed gradient as the compressed differential, the construction of the differential no longer depends on the model states from iterations $t$ or tnt-n, eliminating prior data dependencies for DC calculation. As illustrated in Figure 3(b), together these factors ensure that gradient reuse introduces no data dependency from checkpointing to training, thereby decoupling the two processes and allowing them to execute fully in parallel.
Figure 3: Data dependency analysis.
IV Design
Figure 4: Architecture of LowDiff
We design LowDiff based on the main idea in §III-C with the following design goals.
  • Lowering the costs of DC. LowDiff reduces or even eliminates the computation and transmission costs of compressed differential checkpoints by reusing compressed gradients (§IV-A).
  • Enhancing DC performance. LowDiff accelerates the differential checkpointing under high frequency via batched gradient writing optimization (§IV-B).
  • Minimizing the wasted time. LowDiff minimizes the wasted time by analytically tuning the full checkpointing frequency and gradient batching size (§IV-C).
Figure 4 shows the architecture of LowDiff. Each worker has two processes: Training and Checkpointing, connected via Reusing Queue for gradient reuse. Training first generates the local gradient, compresses it in Compressor, and synchronizes it across workers via Communicator. The synchronized compressed gradient $\widetilde{G}_{t}$ is put into Reusing Queue via zero-copy technique for reuse in checkpointing and is also decompressed by Decompressor to update the model state. Training checkpoints the model state $M_{t}$ regularly as full checkpoints $C^{F}_{t}$. Checkpointing snapshots compressed gradients from Reusing Queue to form differential checkpoints $C^{D}_{t}$, which Batched Writing module caches in CPU memory and aggregates into a single batched differential checkpoint $C^{B}_{t}$. Both the full checkpoint frequency and batching size are dynamically tuned by Optimal Configuration module. Finally, the batched differential and full checkpoints are persisted to local or remote storage.
IV-A Reusing Compressed Gradients
Design requirements. The reuse of the compressed gradients to act as the compressed differentials imposes two main system-level requirements: Requirement 1: Ensuring sequential order for reusing gradients. Since the compressed gradients are generated sequentially during training and differential checkpoints also need to capture model state changes sequentially (see Equation (2)), it is necessary to preserve the sequence of the compressed gradients to ensure that they act as the differentials in the correct sequential order. Requirement 2: Facilitating transmission between training and checkpointing. To avoid performance degradation, the reused gradients need to be transmitted between training and checkpointing with low communication overhead.
Reusing queue structure. To meet Requirement 1, we design a queue-based data structure, referred to as the compressed gradient Reusing Queue in Figure 4. Specifically, the first-in-first-out (FIFO) property of the queue naturally aligns with the iterative nature of training and checkpointing, ensuring that gradients are reused in the correct order. In addition, the implementation of the queue structure in PyTorch also helps meet Requirement 2, which will be specified below.
Reusing with zero-copy optimization. To meet Requirement 2, we use the CUDA Inter-Process Communication (IPC) queue in PyTorch to implement a zero-copy mechanism. This allows GPU memory to be shared between training and checkpointing without extra copying, significantly reducing communication overhead. Specifically, the queue serves as a high-level data structure for reusing compressed gradients, but only transmits the memory handle. It relies on the CUDA IPC library to share the gradient’s GPU memory handle across processes, avoiding additional I/O overhead from memory copying. In PyTorch, this can be achieved with torch.multiprocessing.Queue(), where the inter-process communication creates an IPC handle when a tensor is enqueued and reopens it when dequeued.
Design details Algorithm 1 specifies the following details of reusing compressed gradients, including training, checkpointing, and recovery processes.
Training process. After performing a backward pass and generating the gradient $G_{i,t}$ of worker $i$ at the $t$-th iteration (Lines 1-3), the training process compresses $G_{i,t}$ into $\widetilde{G}_{i,t}$ for communication (Line 4). Then, the compressed gradient is synchronized across all workers using a collective communication primitive (e.g., Allgather or Allreduce) (Line 5). Following synchronization, the training process puts the compressed gradient $\widetilde{G}_{t}$ into Reusing Queue $Q$ (Line 6), which will be transmitted to the checkpointing process with zero-copy. Finally, the training process decompresses the gradient (Line 6), updates the model state based on the gradient (Line 8), and proceeds to the next iteration.
Checkpointing process. The checkpointing process repeatedly retrieves the compressed gradients from Reusing Queue $Q$ to CPU memory (Line 11). Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §IV-B). After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12-14) to storage. The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 15).
Recovery process. During recovery, the recovery process first loads the latest full checkpoint $C^{F}_{t}$ at the $t$-th iteration to restore the model state $M_{t}$ (Line 17). Then, the compressed gradient $\widetilde{G}_{j}$ is retrieved from the differential checkpoint $C^{D}_{j}$ at the $j$-th iteration (Line 20) and decompressed into gradient $G_{j}$ (Line 21). Subsequently, the model state is restored from $M_{j}$ to $M_{j+1}$ at the $(j)$-th iteration by the optimizer and gradient (Line 21). By repeating this process for $n$ iterations (Lines 18-23), where compressed gradients from differential checkpoints are successively loaded and applied, the model state is finally restored to $M_{n}$ (Line 24), thus recovering the training from failures.
Algorithm 1 Reusing in LowDiff
Model state $M_{t}$at iteration $t$, Gradient $G_{i,t}$at worker $i$, Reusing queue $Q$, Full checkpoint $C^{F}_{t}$, Differential checkpoint $C^{D}_{n}$at iteration $n$, Compressor CompComp, Decompressor $Comp^{-1}$. Training Process $t$
[…]
IV-B Batched Gradient Writing Optimization
Limitations of LowDiff. We find that LowDiff in §IV-A still has two limitations: Limitation 1: Many writes due to frequent DC. The high frequency leads to numerous differential checkpoints, which incur significant write overhead. Limitation 2: Increased GPU memory consumption from frequent gradient reuse. The high frequency of gradient reuse results in many compressed gradients being retained in GPU memory, waiting to be written to storage.
Batched Write. To address Limitation 1, we can batch multiple writes of compressed gradients (reused as differential checkpoints) into a single write operation, with the help of the widely-used gradient accumulation technique [30, 22, 2] where the gradients of the same shape and size can be accumulated. Note that the above batching scheme, which benefits DC writes, cannot work for traditional checkpointing that only saves full checkpoints, since if a new full checkpoint comes, the previous full checkpoints become obsolete and can be discarded without the need to accumulate them.
Offloading batching to CPU. To address Limitation 2, we adopt a classic GPU-CPU offloading scheme [42] that offloads the batching operation to the CPU and caches the compressed gradients in CPU memory, thus reducing the GPU memory cost caused by frequent checkpointing. Note that the batching operation can be easily handled by the CPU since the operation mainly involves the addition of compressed gradients.
Design details To solve Limitation 1 & 2, we design a batched gradient writing optimization, which offloads multiple compressed gradients from GPU to CPU memory, and then aggregates them into batched writes to storage. As shown in Figure 5, we specify the batched gradient writing optimization of LowDiff in three steps:
Step ①: Offloading gradients to CPU memory. After the training process transmits compressed gradients $\widetilde{G}_{t}$, the checkpointing process retrieves the GPU memory handles and offloads the gradients to CPU memory. After offloading, the checkpointing process closes the handle of $\widetilde{G}_{t}$ and frees the corresponding GPU memory.
Step ②: Batching gradients in buffer. The compressed gradients, which are reused to act as differential checkpoints $C^{D}_{t}$, are buffered in CPU memory, awaiting batching. Once the batching size is reached (e.g., the size is four in Fig 5), the checkpointing process groups the buffered differential checkpoints $C^{D}_{t}$, $C^{D}_{t-1}$, $C^{D}_{t-2}$, and $C^{D}_{t-3}$ into a batch checkpoint $C^{B}_{t}$.
Step ③: Writing batched checkpoint to storage. The checkpointing process extracts the batched differential checkpoint $C^{B}_{t}$ from the CPU buffer and writes it to storage in a single I/O operation.
Figure 5: Batched gradient write optimization
IV-C Optimizing Checkpointing Configuration
Insight into configuring LowDiff. Note that recent studies on optimizing the checkpointing performance only focus on configuring the checkpointing frequency (e.g., [19]). However, we observe an insight that LowDiff’s checkpointing performance will also rely on the batching size, i.e., the amount of the batched gradients every batching, introduced in §IV-B. To verify the insight, we take another measurement analysis with the same setting as that in §III-A. Table I shows the results of the wasted time, which evaluates the checkpointing performance (see §II-B), under different batching sizes and checkpointing frequencies for full checkpoint (called full checkpoint frequency, FCF for short). Note that all the results are normalized by comparing them to the lowest wasted time.
We first see that when FCF is too low ($FCF=10$) or too high ($FCF=100$), the wasted time will be increased, since a low FCF will incur high recovery time while a high FCF will incur high steady state overhead (GPU time), Thus, we need to tune FCF to minimize the wasted time, which has been found by Microsoft [19].
More importantly, under the same FCF (i.e., in the same row), we also need to tune the batching size (BS for short) for minimizing the wasted time, For example, when $FCF=10$ or $20$, the wasted time is minimal when $BS=2$; when $FCF=50$ or $100$, the wasted time is minimal when $BS=3$. The reason is that a low BS still incurs many writes due to insufficient batching, while a high BS makes one failure cause many gradients within the batch to be lost.
Therefore, we need to tune both FCF and BS to minimize LowDiff’s wasted time. For example, the wasted time in Table I achieves its minimum when $FCF=20$ and $BS=2$.
FCF BS 1 2 3 4 5 6
10 1.258 1.226 1.279 1.330 1.408 1.513
20 1.052 1.000 1.031 1.085 1.149 1.202
50 1.394 1.290 1.255 1.312 1.387 1.481
100 1.916 1.783 1.712 1.752 1.869 1.952
TABLE I: Normalized wasted time with different full checkpoint frequency (FCF) and batching size (BS).
Configuration Modeling. To find the configuration of FCF and BS for minimizing the wasted time, we model the configuration by constructing the expression of the wasted time with respect to both FCF ($f$) and BS ($b$) variables. We also list the constant parameters of distributed training with differential checkpointing: the number of GPUs $N$, the mean time between failures $M$, the checkpointing write bandwidth $W$, a full checkpoint size $S$, the total run-time for a training job $T$, the time to load a full checkpoint $R_{F}$, and the time to merge a differential checkpoint $R_{D}$ with the full one.
The following expression can be obtained based on the two variables and the constant system parameters above:
  • The number of failures for a training job = $\frac{T}{M}$.
  • The time of writing a full checkpoint = $\frac{S}{W}$.
  • The number of full checkpoints during training = $f\times T$.
  • The wasted work which is re-processed upon the last checkpoint across all GPUs = $N\times\frac{T}{M}\times\frac{b}{2}$. Note that half-batched checkpoints might be lost on average in the event of failures.
  • The number of merging operations on average = $\frac{1}{2}(\frac{1}{f}\times\frac{1}{b}-1)$. Note that the number of batched DC equals dividing the full checkpoint interval by the batching size.
  • The recovery time for all nodes to load checkpoints = $N\times\frac{T}{M}\times(R_{F}+R_{D}\times\frac{1}{2}(\frac{1}{f}\times\frac{1}{b}-1))$, including loading a full checkpoint and merging differential checkpoints.
  • Total recovery overhead due to failures across all GPUs= $N\times\frac{T}{M}\times((R_{F}+R_{D}\times\frac{1}{2}(\frac{1}{f}\times\frac{1}{b}-1))+\frac{b}{2})$, including loading checkpoints and re-processing the wasted work.
  • Total steady state overhead for checkpointing across all GPUs = $N\times\frac{S}{W}\times f\times T$.
The wasted time $T_{wasted}$ can be calculated by adding recovery overhead and steady state overhead (see §II-B):
$T_{wasted}$ $=\frac{NT}{M}(\frac{b}{2}+(R_{F}+\frac{R_{D}}{2}(\frac{1}{f\times b}-1)))+\frac{NT\times S\times f}{W}.$ (3)
To minimize $T_{wasted}$ on two variables $f$ and $b$, we take the first-order partial derivatives of $T_{wasted}$ with respect to $f$ and $b$.
$\begin{cases}\frac{\partial T_{wasted}}{\partial f}=\frac{NST}{W}-\frac{NT\times R_{D}}{2f^{2}\times M\times b}=0.\\ \frac{\partial T_{wasted}}{\partial b}=\frac{1}{2}-\frac{R_{D}}{2b^{2}\times f}=0.\end{cases}$ (4)
Equation (4) can be reduced to obtain the optimal values of the full checkpointing frequency ($f^{*}$) and batching size ($b^{*}$) as follows.
$(f^{*},b^{*})=(\sqrt[3]{\frac{R_{D}W^{2}}{4S^{2}M^{2}}},\sqrt[3]{\frac{2SR_{D}M}{W}})$ (5)
We will use $(f^{*},b^{*})$ as the optimal configuration in LowDiff.
V enhancement
Figure 6: Layer-wise gradient reuse.
Issue of LowDiff without gradient compression. The current design of LowDiff relies on gradient compression to enable efficient gradient reuse. When compression is disabled, we can still enforce a lightweight, reuse-only compression to preserve the bandwidth reduction in Finding 2, but the computational savings of Finding 1 no longer apply. In this case, the differential checkpoints become significantly larger, slowing down their snapshotting and persistence accordingly. As a result, checkpointing competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall introduced by gradient reuse, snapshotting, and persistence of DC under such non-compressed scenarios. We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
V-A Layer-wise gradient reusing and snapshotting.
A key challenge in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of blocking. To accelerate such large-scale snapshotting, we aim to design a fine-grained strategy that better overlaps gradient transmission with training.
As shown in Figure 6, gradients are produced incrementally in the backward pass, following the reverse layer order during training. Instead of waiting until the entire backward pass finishes to snapshot all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [10], PyTorch DDP [28], and Horovod [46]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced. Inspired by this execution model, we design a fine-grained reuse strategy: immediately snapshot each layer’s gradient to host memory upon computation and stream it into a differential checkpoint assembler. By pipelining backpropagation, communication, and gradient reuse at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
Figure 7: Batching strategies of LowDiff and LowDiff+.
V-B CPU-based asynchronous persistence strategy.
The layer-wise strategy addresses the snapshotting issue in LowDiff with compressed gradients. Another challenge lies in the persistence of both full and differential checkpoints. The batching strategy described in §IV-B accumulates small compressed updates on the CPU side before writing them out. It also periodically persists the full model state, which works effectively when update sizes are small and sparse. In the non-compression setting, differential checkpoints are significantly larger, and naively batching them demands a sustained write throughput that exceeds the available storage bandwidth, making it essential to design a more efficient checkpoint merging strategy that better lowers the writing data volume.
To address this, we observe that in the non-compression setting, gradients do not need to be persisted separately. Instead, they can be directly fused with the full model state inside CPU memory, enabling the CPU to maintain an always-up-to-date replica of the GPU model. Once gradients are merged into the CPU-resident model state, persistence only needs to occur from the CPU model to storage. By fusing gradients into the CPU-resident model state instead of persisting them separately, redundant writes of differential checkpoints are avoided, which directly reduces the overall storage volume. The different batching strategies of LowDiff and LowDiff+ are shown in 7. This strategy decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
V-C Design of LowDiff+
Design details. Based on Insight 1 and 2, we design LowDiff+ without gradient compression. The architecture of LowDiff+ is shown in Figure 8, and its detailed algorithm is presented in Algorithm 2. In the training process, the model first performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer as shown in Figure 6 (Line 5). The layer-wise gradient communication is then handled by a communication thread pool $P_{g}$, where the gradient synchronization is assigned to multiple communication threads (Line 15). Once the communication in a thread finishes, the corresponding gradients are forwarded via the reusing queue to the CPU-side checkpointing process (Line 16), while the training process waits for the full gradient $G_{t}$ (i.e., wait for handle set $H_{g}$) (Line 6) to update the GPU model $M^{G}_{t}$ (Line 7).
Figure 8: Workflow of LowDiff without gradient compression.
In the checkpointing process, gradients $G^{n}_{t}$ streamed from the GPU are reused layer by layer (Line 9). The offloading thread pool $P_{s}$ handles multiple layers in parallel (Line 10), snapshotting the layer-wise gradients into CPU memory as they become available (Line 19). We initialize a CPU-resident replica $M^{C}_{t}$ of the model parameters, aligned with the GPU state. After snapshotting (Line 11), reused gradients are applied to $M^{C}_{t}$ to keep it synchronized with the GPU model $M^{G}_{t}$ (Line 12). The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [54]. Since such updates do not involve forward or backward computations, their execution on the CPU introduces negligible overhead. Persistence is performed asynchronously by checkpointing the CPU-maintained model replica (Line 13), rather than the raw gradients. The frequency of persistence is decided similarly to Checkfreq [36], which ensures maximal overlapping between the persist and update in CPU. This design both avoids redundant storage of differential checkpoints and completely decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
When failure occurs, prior work [54, 19] categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures. For hardware failures (e.g., GPU device crashes), the distributed training system must replace the failed machine with a new one. In this case, the in-memory training state is lost, so the recovery process must reload model states from persistent storage checkpoints, which can be time-consuming and increase wasted time. In contrast, software failures (e.g., software bugs) only cause the training process to terminate while leaving the CPU memory state of the checkpointing process intact. In such cases, recovery is achieved by reinitializing the training process and restoring the preserved CPU-resident model state to the GPUs, leveraging in-memory checkpoints. This enables fast recovery under software failures, avoiding expensive checkpoint reloads and serving as a key optimization of LowDiff+.
Algorithm 2 Reusing in LowDiff+
[…]
VI Implementation
VI-A Implementation of LowDiff
We implement LowDiff atop PyTorch [40, 39] and DeepSpeed [10, 44], consisting of three main modules: reusing queue (§IV-A), batched gradient writing (§IV-B), and optimal configuration (§IV-C). In addition, a parallel recovery module accelerates recovery from a large number of differential checkpoints at high frequency.
Reusing queue module. Since Python’s GIL prevents true multi-threading, LowDiff uses torch.multiprocessing with the spawn method for process isolation. The compressed gradient Reusing Queue is implemented via torch.multiprocessing.Queue(), which (1) preserves checkpoint order and (2) leverages CUDA IPC for zero-copy GPU memory sharing.
Batched gradient writing module. Differential checkpoints are offloaded from GPU to CPU memory with offload_to_cpu() (non-blocking) and temporarily cached in a CPU buffer. GPU memory handles are released after offloading. Once the batch size is reached, checkpoints are aggregated (tensor addition or dictionary accumulation) and persisted to local or remote storage.
Optimal configuration module. This module tunes full-checkpoint frequency and batch size based on Eq. (5). Starting from a default setup, it adapts to runtime metrics (failure rate, bandwidth) using stepwise adjustments to balance steady-state overhead and recovery cost, achieving efficient fault-tolerant checkpointing.
Parallel recovery module.Traditional recovery methods for differential checkpointing [15] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (2)), whose the number of merge operations is $n$ with given $n$ differential checkpoints. Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints. Specifically, as shown in figure 9, it merges the earliest differential checkpoint $C_{1}^{D}$ with the full checkpoint $C_{t}^{F}$ while concurrently merging the subsequent differential checkpoints in pairs (i.e., $C_{2}^{D}$ and $C_{3}^{D}$, $C_{4}^{D}$ and $C_{5}^{D}$). This parallelization reduces the number of merge operations between differential checkpoint and full checkpoint from $n$ to $\log n$. In Figure 9, the number of merges between differential checkpoint and full checkpoint is 2 (the serial process has $n=5$ merges). Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
VI-B Implementation of LowDiff+
LowDiff+ is also implemented atop PyTorch [40, 39] and DeepSpeed [10, 44]. Since the checkpointing process is spawned from the training process, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning. For gradient communication, we leverage Python’s concurrent.futures library and employ a ThreadPoolExecutor to assign each layer-wise gradient to a communication thread, ensuring parallel transmission. The snapshotting thread pool is implemented in the same manner, handling the writing of gradients into CPU memory.
To maintain the CPU-side model replica, we configure PyTorch intra-op parallelism via torch.set_num_threads, setting the number of threads to match the available CPU cores in order to maximize parallel efficiency. Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation. Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [36], thereby maximizing checkpointing throughput without interfering with GPU training.
Figure 9: Parallel Fast Recovery
VII Evaluation
In this section, we evaluate the efficiency of LowDiff to answer the following questions:
  • How fast do LowDiff train under high checkpointing frequency compared to other methods? (Exp. 1)
  • How fast do LowDiff+ train under non-compression scenarios? (Exp. 2)
  • Can LowDiff and LowDiff+ achieve a shorter wasted time? (Exp. 3)
  • Does LowDiff and LowDiff+ enable higher checkpointing frequency under bounded training speed? (Exp. 4)
  • Does LowDiff and LowDiff+ accelerate the recovery process? (Exp. 5)
  • Can LowDiff’s batched writes and offloaded batching reduce the checkpointing time and GPU memory cost? (Exp. 6)
  • Can LowDiff reduce the storage overhead? (Exp. 7)
  • Does LowDiff achieve frequent checkpointing under different compression ratios? (Exp. 8)
  • Can LowDiff and LowDiff+ maintain high training efficiency under scenarios of more frequent failures and more GPUs? (Exp. 9-10)
GPU Type GPU Mem CPU Type
NVIDIA A100 80GB Intel Xeon 8352V
NVIDIA V100S 32GB Intel Xeon 4214
(a) GPU and CPU configurations of servers
Models Datasets Params
ResNet-50 Cifar-100 25.6M
ResNet-101 ImageNet 44.5M
VGG-16 Cifar-100 138.8M
VGG-19 ImageNet 143.7M
BERT-B SQuAD 110M
BERT-L SQuAD 334M
GPT2-S WikiText-2 117M
GPT2-L WikiText-103 762M
(b) Models and datasets used for evaluation
TABLE II: Experimental setup
VII-A Experimental Setup
Servers. We assess the performance of LowDiff on two generations of NVIDIA GPUs, the A100 and V100S, with different CPU configurations. The specific server configurations are detailed in Table II(a). Each server is equipped with 4 GPUs, 512GB of system memory, and a 4TB Samsung SSD. Within a server, GPUs are interconnected via NVLink, while cross-server communication is performed over a 25Gbps Mellanox ConnectX-5 InfiniBand network. The A100 servers are equipped with PCIe Gen 4 interfaces, whereas the V100S servers use PCIe Gen 3 interfaces. All servers operate on Ubuntu 22.04 and are equipped with the following software libraries: Deepspeed-0.16.4, CUDA-12.4, NCCL-2.23.4, OpenMPI-4.0.5, Python-3.10.15, and PyTorch-2.6.0.
Workloads. We evaluate widely-used DNN models on tasks of image classification and natural language processing (NLP), including ResNet [20], VGG [49], Bert [14], and GPT-2 [41]. The image classification models are evaluated on the CIFAR-100 [25] and ImageNet [12] datasets, and NLP models are evaluated on the SQuAD [43] and WikiText [33] datasets. The detailed configurations are shown in Table II(b). We use Adam [24] as the default optimizer. The training process uses sparsification compression with a common compress ratio $\rho$ of 0.01 [3, 29, 18, 35, 7, 45]. We use default batching sizes and learning rates according to the original papers or standard practices. We also implement the LowDiff system on the VGG-16 model with pipeline parallelism provided in the DeepSpeedExamples [11].
Baselines. We evaluate two state-of-the-art checkpointing strategies: CheckFreq [36] and Gemini [54], as described in §II-B. The experiments follow the default configurations according to their original papers. Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [15] does in recommendation systems.
VII-B Performance Evaluation
Figure 10: (Exp. 1) Training time of different models with different checkpointing strategies.
Experiment 1 (Training time of LowDiff). We evaluate the training time of LowDiff under different checkpointing strategies on the previously introduced DNN models, with the checkpointing frequency set to once per iteration. The experiment is conducted in the presence of gradient compression. For each model, we run 1,000 training iterations on A100 servers and record the total training time. The W/O CKPT setting represents training without checkpointing, which serves as the upper bound of achievable training speed.
Figure 10 shows that the training time of LowDiff on all seven training tasks is close to that of W/O CKPT and significantly smaller than those of other solutions. Compared to W/O CKPT, LowDiff’s training time increases by only 2.4%-3.1%, while the other checkpointing methods’ training time increases by 8.1%-891%.
Specifically, we elaborate on LowDiff’s benefits over the other methods and analyze the reasons behind these benefits:
  • Compared to CheckFreq and Gemini in GPT2-S, LowDiff reduces the training time by 68.2% and 46.1%, respectively. We see that LowDiff significantly outperforms them under high frequency since LowDiff uses differential checkpointing with compression (compression ratio=0.01) to largely reduce the checkpoint sizes. Further, Gemini improves its training performance over CheckFreq by checkpointing to CPU memory, but the improvement is limited by its large checkpoint size.
  • LowDiff’s performance improvement is more pronounced for larger models like BERT-L and GPT2-L. Compared to CheckFreq and Gemini in GPT2-L, LowDiff reduces the training time by 89.2% and 59.2%, respectively. This is because, as the model size increases, the advantage of the reduced checkpoint size of LowDiff becomes more dominant on the overall training time.
  • Compared to Naïve DC in BERT-B, LowDiff reduces the training time by 60.5% due to its unique gradient reusing scheme, which reduces and even eliminates compression and transmission cost of Naïve DC, confirming the efficiency of our reusing idea in §III-C. It should be noted that, in comparison with CheckFreq and Gemini, Naïve DC fails to demonstrate the advantage of its reduced checkpoint size due to its high costs.
  • LowDiff’s performance improvement remains significant in pipeline parallelism. Compared to Naïve DC, CheckFreq, and Gemini in VGG16 with pipeline parallelism, LowDiff reduces the training time by 70.8%, 60.9%, and 36.9%, respectively. The reason is that LowDiff’s efficient high-frequency DC relies on the core technique that reuses the compressed gradients, which are computed during training and still exist in pipeline-parallelism-based training systems. Thus, reusing compressed gradients and pipeline parallelism can be combined for training efficiency. Our future work will explore the efficiency of such a combination.
Figure 11: (Exp. 2) Training time without compression.
Experiment 2 (Training time of LowDiff+). To isolate the impact of gradient compression, we further evaluate the training time of LowDiff+ in a setting without gradient compression. All other experimental configurations, including the checkpointing frequency (per iteration), the selected DNN models, the number of training iterations (1,000), and the hardware environment (A100 servers), remain identical to Exp. 1.
Figure 11 shows that the training time of LowDiff+ across all tasks exhibits a slight increase compared to the W/O CKPT baseline, yet it remains significantly lower than that of other checkpointing solutions. Specifically, LowDiff+’s training time increases by 8.2%–10.1% relative to W/O CKPT, which is slightly higher than the increase observed in the setting without gradient compression. This overhead is primarily due to the frequent and large-volume gradient transfers between GPU and CPU memory, which occupy PCIe bandwidth and consequently reduce training performance.
In the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods. For the GPT2-L training task, it reduces the training time by 51.8% compared to Gemini and by 81.7% compared to CheckFreq, demonstrating its superior efficiency in settings where frequent non-compression checkpoints are required.
Figure 12: (Exp. 3) Wasted time under different MTBFs.
Experiment 3 (Wasted time). We evaluate the wasted time of various checkpointing methods in the event of failures. Failures were simulated during training by adhering to a fixed Mean Time Between Failures (MTBF) metric. Using three MTBF values at 0.5, 1, and 2 [19], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S. Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §IV-C. For LowDiff+, we evaluate its wasted time under both software and hardware failures (LowDiff+ (S) and LowDiff+ (H), respectively).
Figure 12 shows that the wasted time gap between LowDiff and other checkpointing methods keeps increasing as failures become more frequent, while LowDiff maintains the lowest wasted time. When the MTBF decreases from 2 to 0.5 (lower MTBF means more frequent failures), the gap between LowDiff and Gemini increases from 0.061h to 0.145h. This is because LowDiff tunes the full checkpointing frequency and batching size to approach the optimal configuration of checkpointing, while other methods’ wasted time is greatly influenced by the different parameters. The results confirm the efficiency of LowDiff in the event of failures.
For LowDiff+, the wasted time due to software failures is 3.7%–5.1% lower than that of LowDiff. This improvement comes from LowDiff+ ’s in-memory checkpointing, which enables fast recovery using the full model state stored in CPU memory. The detailed recovery times of LowDiff+ (S) are presented in Exp. 5. In contrast, the wasted time from hardware failures is slightly higher than that of LowDiff, due to LowDiff+ ’s lower checkpointing frequency, but still lower than that of CheckFreq and Gemini. Details of the checkpointing frequency for LowDiff+ (S) and LowDiff+ (H) are presented in Exp. 6.
Figure 13: (Exp. 4) Maximum checkpointing frequency.
Experiment 4 (Checkpointing frequency). To evaluate the maximum checkpointing frequency that each method can achieve under bounded training speed (set as 3.5%, as the degraded speed of model training is required to be less than 3.5% in Microsoft’s study [36]), we evaluate the highest checkpointing frequency of different methods. We denote LowDiff+ ’s in-memory checkpointing as LowDiff+ (S) and its persistence as LowDiff+ (P).
Figure 13 shows that compared to other methods, LowDiff can achieve the checkpointing frequency every iteration (i.e., the checkpointing frequency equals one) for different widely-used models under the bounded training speed.
LowDiff+ (S) also performs per-iteration in-memory checkpointing, enabled by its efficient layer-wise gradient reuse and snapshotting approach. LowDiff+ (P) achieves per-iteration persistence frequency in smaller models such as ResNet-101, but increases to 3 iterations for larger models like GPT2-L due to the limited bandwidth of SSDs. Nevertheless, its checkpointing intervals remain the shortest compared to Naïve DC, CheckFreq, and Gemini.
Intervals of Naïve DC grow significantly with model size, increasing from every 2 iterations to every 8. Gemini only supports per-iteration checkpointing on ResNet-101, and its interval increases to every 4 iterations for GPT2-L and BERT-L. CheckFreq consistently maintains an interval of 10 iterations. These results validate our motivation that LowDiff can support high-frequency checkpointing without stalling training progress.
Figure 14: (Exp. 5) Recovery time of different methods.
Experiment 5 (Recovery time). We evaluate the recovery time of different recovery strategies under different full checkpointing frequency in GPT2-S. We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [39]), Naïve DC, LowDiff with parallel recovery (see §VI), and LowDiff+ from software failures (LowDiff+ (S) for short). Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
[…]
(a) Average checkpointing time under different batch sizes
(b) GPU memory consumption w/o and w/ offloaded batching
Figure 15: (Exp. 6) CKPT time and GPU memory reduction.
Experiment 6 (Checkpointing time reduction and GPU memory reduction). To evaluate the reduction in average checkpointing time achieved by the batched write optimization proposed in §IV-B, we conduct experiments on various models with different batching sizes. Figure 15a shows that the batched write strategy significantly reduces the average time required to write differential checkpoints, with a maximum reduction of $30.9\%$ when the batching size is 20 during GPT2-S training.
The batching strategy effectively accelerates the checkpoint writing process. After tuning the batching frequency with the configuration strategies discussed in §IV-A, the optimization can adapt to different compression ratios and minimize the impact on training performance.
Additionally, we also evaluate the reduction of GPU memory achieved by the batching offloaded in CPU §IV-B. Figure 15b shows that the GPU memory usage will increase by $10\%-12\%$ without offloaded batching, especially for GPT2-L, potentially leading to memory exhaustion. In contrast, the GPU memory usage will return to the original state with the help of offloaded batching.
Model Full CKPT Naïve DC LowDiff
ResNet-101 511M 346M 34M
VGG-19 1.7G 1.13G 109M
BERT-B 1.3G 930M 82M
BERT-L 3.8G 2.55G 239M
GPT2-S 1.4G 946M 92M
GPT2-L 8.7G 5.7G 541M
TABLE III: (Exp. 7) Storage overhead of checkpoints.
Experiment 7 (Storage overhead). We evaluate the storage overhead of checkpoints for various checkpointing methods and various DNN models, where CheckFreq and Gemini only have full checkpoints (called Full CKPT), Naïve DC and LowDiff have full and differential checkpoints. As shown in Table III, compared to full checkpoints, Naïve DC reduces the storage overhead by 34.4% due to the compression technique used in the model parameters.
Compared to Naïve DC, LowDiff further reduces the storage overhead by 90.5%. Because LowDiff utilizes compressed gradients to construct differentials for both model parameters and optimizer parameters (Finding 2 in §III-B), while Naïve DC (Check-N-Run [15]) does not apply sparsification compression to the optimizer parameters within differential checkpoints, resulting in the optimizer occupying the majority of the checkpoint space. The result ensures our motivation that LowDiff can significantly reduce the checkpoint size and thus lower the storage cost.
Figure 16: (Exp. 8) Impact of Compression Ratio $\rho$.
Experiment 8 (Impact of compression ratio $\rho$). We evaluate the impact of different compression ratios $\rho$ on the checkpoint frequency of GPT2-L and GPT2-S training. The compression ratio $\rho$ varies from 0.1 to 0.001, consistent with the ranges commonly used in prior work [29, 35].
As shown in Figure 16, LowDiff enables per-iteration checkpointing for GPT2-S training across the entire range of $\rho\in[0.001,0.1]$ due to its smaller enough gradient size for checkpointing to be overlapped with training within one iteration. For GPT2-L, LowDiff supports per-iteration checkpointing when $\rho\in[0.001,0.075]$, while at $\rho=0.1$, the checkpoint frequency increases to 2 iterations because the increasing gradient size makes the checkpointing take two iterations to be overlapped with training. The result shows LowDiff ensures frequent checkpointing (intervals ¡ 3 iterations) under the common compression ratio.
VII-C System Scalability
We test the scalability of LowDiff and LowDiff+ on V100 servers in scenarios with frequent failures and more GPUs to demonstrate the robustness and efficiency of LowDiff and LowDiff+.
Experiment 9 (Scaling to frequent failures). We conduct experiments on the performance of LowDiff and LowDiff+ in extreme scenarios characterized by more frequent failures of the training process with MTBF ranging from 0.1 to 5 hours. Using the effective training time ratio metric as defined in Gemini [54], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
Figure 17 demonstrates that LowDiff consistently sustains a higher effective training time ratio across all tested scenarios, especially under environments with frequent failures. Although LowDiff+ achieves slightly lower ratios than LowDiff, it still outperforms the other methods. For example, when the MTBF is set to 0.3 hours, LowDiff reaches an effective training time ratio of 92%, while LowDiff+ achieves 86%, compared to only 81% for Gemini and 76% for Checkfreq. These results highlight the robustness and efficiency of both LowDiff and LowDiff+ in handling frequent failures during training.
Experiment 10 (Scaling to more GPUs). We evaluate the training performance by varying the number of GPUs used in the training process, using the same effective training time ratio metric as in Experiment 7. As the number of GPUs increases, the probability of failures in the cluster also rises [19]. We conduct experiments with 8, 16, 32, and 64 GPUs, respectively, and measure the effective training time ratio of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
Figure 18 shows that, as the number of GPUs increases, the effective training time ratio decreases; however, LowDiff still maintains a 98% proportion of the effective training time ratio, and LowDiff+ achieves 96%, while the metric for the other method declines rapidly and can only reach around 90%. This highlights the strong potential of LowDiff and LowDiff+ in large-scale training.
Figure 17: (Exp. 9) Training under frequent failures.
Figure 18: (Exp. 10) Training with different number of GPUs.
VIII Conclusion and Future Work
LowDiff is an efficient frequent checkpointing framework that addresses the challenges of the high costs of differential checkpointing by reusing the compressed gradients generated during training. LowDiff also designs a batched gradient writing optimization scheme and an optimal configuration tuning scheme. To enhance the gradient reuse strategy in non-compression scenarios, we design LowDiff+. LowDiff+ uses a layer-wise gradient reusing and snapshotting strategy and a CPU-based synchronous persistence strategy, thus enabling checkpointing to decouple from training. Our experiments on a variety of models and GPU generations demonstrate the efficiency of LowDiff and LowDiff+ in terms of the training time, wasted time, checkpointing time, and recovery time under the high-frequency checkpointing scenarios.

### v2

@@
-LowDiff: Efficient Frequent Checkpointing via Low-Cost Differential for High-Performance Distributed Training Systems
+Optimizing Frequent Checkpointing via Low-Cost Differential for Distributed Training Systems
 Abstract
@@
 Recently, differential checkpointing has been proposed to reduce costs, but it is limited to recommendation systems, so its application to general distributed training systems remains unexplored.
-This paper proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost.
+We proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost.
 Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently.
 It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance.
-We further enhance LowDiff with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression.
+In non-compression scenario, We further proposes LowDiff+ with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression.
 Experiments on various workloads show that LowDiff can achieve checkpointing frequency up to per iteration with less than 3.1% runtime overhead.
 I Introduction
-Deep learning has demonstrated remarkable capabilities in the fields of computer vision [49, 20], natural language processing [14, 6] and so on.
-Recently, the advent of large models like ChatGPT [37] and GPT-4 [38] has captivated both academic and industrial attention.
-To efficiently train the large models, distributed training systems [27, 26, 46, 44] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations.
-Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [55].
-For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [23]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [34].
+Deep learning has demonstrated remarkable capabilities in the fields of computer vision [simonyan2014very, he2016deep], natural language processing [devlin2018bert, brown2020language] and so on.
+Recently, the advent of large models like ChatGPT [chatgpt] and GPT-4 [openai2024gpt4] has captivated both academic and industrial attention.
+To efficiently train the large models, distributed training systems [li2013parameter, li2014communication, sergeev2018horovod, rasley2020deepspeed] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations.
+Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [zhang2020empirical].
+For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [jeon2019analysis]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [OPT].
 To handle failures, checkpointing, which regularly writes the model state to storage (e.g., disks) for failure recovery, can confine the potential loss of the training progress to the interval between consecutive checkpoints, only requiring the recovery of the remaining progress.
-Traditional checkpointing saves the model states (including the model parameters and optimizer parameters) at epoch boundaries, but an epoch often runs for hours, so the epoch-level checkpointing often results in hours of GPU computation wasted on average for every failure recovery [36].
-Consequently, state-of-the-art checkpointing studies on fast recovery (e.g., Checkfreq [36] and Gemini [54]) aim to increase the frequency of checkpointing from epochs to iterations (we call frequent checkpointing).
+Traditional checkpointing saves the model states (including the model parameters and optimizer parameters) at epoch boundaries, but an epoch often runs for hours, so the epoch-level checkpointing often results in hours of GPU computation wasted on average for every failure recovery [mohan2021checkfreq].
+Consequently, state-of-the-art checkpointing studies on fast recovery (e.g., Checkfreq [mohan2021checkfreq] and Gemini [wang2023gemini]) aim to increase the frequency of checkpointing from epochs to iterations (we call frequent checkpointing).
 Nevertheless, the frequent checkpointing generates a considerable number of checkpoints, whose creation and transmission will incur substantial computational and communication costs, and thus hinder the training process significantly.
-Recent studies on reducing such costs (e.g., Check-N-Run [15]) introduce differential checkpointing (DC for short), which only tracks and checkpoints the modified part of the model to reduce the checkpoint size, into recommendation systems by leveraging its unique sparse features.
-However, when deploying DC with high frequency in general distributed training systems, we observe two challenges arising: 1) without the help of sparse features, frequent DC incurs significant computation cost required for compressing the original differentials into smaller ones (i.e., compressed differentials), thus hindering the training process, and 2) frequent DC causes significant training stalls, as frequent checkpoint writes incur substantial transmissions which can not be overlapped with training, thus degrading the training performance (see §III-A for details).
-Fortunately, we also have two findings from the gradient compression techniques [7, 18, 35, 9, 50] widely used in distributed training systems for communication efficiency: 1) the compressed gradients can be reused to construct the compressed differentials for general deep-learning models.
-2) Using compressed gradients eliminates the data dependency between training and checkpointing, enabling them to run in parallel.
+Recent studies on reducing such costs (e.g., Check-N-Run [eisenman2022check]) introduce differential checkpointing (DC for short), which only tracks and checkpoints the modified part of the model to reduce the checkpoint size, into recommendation systems by leveraging its unique sparse features.
+However, when deploying DC with high frequency in general distributed training systems, we observe two challenges arising: (1) without the help of sparse features, frequent DC incurs significant computation cost required for compressing the original differentials into smaller ones (i.e., compressed differentials), thus hindering the training process, and (2) frequent DC causes significant training stalls, as frequent checkpoint writes incur substantial transmissions which can not be overlapped with training, thus degrading the training performance (see §III-A for details).
+Fortunately, we also have two findings from the gradient compression techniques [chen2023bbtopk, fei2021efficient, ming2024adtopk, chen2021cswap, song2023optimus] widely used in distributed training systems for communication efficiency: (1) the compressed gradients can be reused to construct the compressed differentials for general deep-learning models.
+(2) Using compressed gradients eliminates the data dependency between training and checkpointing, enabling them to run in parallel.
 Moreover, the size of compressed gradients is significantly smaller than that of compressed differentials (see §III-B for details).
@@
 Our motivation is that the compressed gradients can act as the compressed differentials with no DC computation cost and with a lower DC transmission cost.
-In this paper, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials).
+Based on the findings, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials).
 Further, to enhance LowDiff, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
@@
 Motivated by the two corresponding findings of compressed gradients, we propose LowDiff, which reuses compressed gradients to act as the compressed differential checkpoints, thereby lowering the cost (§III).
+  • We perform an analysis of LowDiff, showing that gradient reuse eliminates the data dependency and enables parallel execution between checkpointing and training.
+Through measurement, we further demonstrate that checkpointing can be overlapped with the training iteration, allowing checkpointing to execute concurrently with training (§IV).
   • We design LowDiff, which realizes the reuse of compressed gradients via a queue structure and facilitates zero-copy transmission between the training and checkpointing processes.
 Atop LowDiff, we propose a batched gradient writing optimization by merging compressed gradients in CPU memory before persisting to storage.
-We also design a checkpointing configuration optimization to minimize the wasted GPU time (§IV).
+We also design a checkpointing configuration optimization to minimize the wasted GPU time (§V).
   • We further propose LowDiff+ for efficient gradient reuse without compression.
 By exploiting the layer-by-layer generation of gradients during training, we design a layer-wise gradient reusing and snapshotting approach.
-We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§V).
-  • We implement LowDiff atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff.
-Experiments on various models show that LowDiff can achieve checkpointing as frequently as once per iteration with less than 3.1% runtime overhead.
-Compared to the state-of-the-art methods, LowDiff can shorten training time by up to $89.2\%$ under high checkpointing frequencies (§VI & §VII).
+We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§VI).
+  • We implement LowDiff and LowDiff+ atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff.
+Experiments on various models show that LowDiff and LowDiff+ can achieve checkpointing as frequently as once per iteration with less than minimal runtime overhead.
+Compared to the state-of-the-art methods, LowDiff and LowDiff+ can shorten training time by up to $89.2\%$ and $81.7\%$ under high checkpointing frequencies (§VII & §VIII).
 II Background and Related Work
@@
 model parameters)
-To accelerate the training process of large models and datasets [56, 28, 46, 44], distributed training systems are commonly adopted to distribute computations and data across multiple GPUs.
+To accelerate the training process of large models and datasets [zhao2023pytorch, li2020pytorch, sergeev2018horovod, rasley2020deepspeed], distributed training systems are commonly adopted to distribute computations and data across multiple GPUs.
 Specifically, distributed training has four steps for the $t$-th iteration:
-  • Forward pass. The model function is applied to the input data set to obtain the prediction.
-  • Backward pass. A loss function is used to calculate the gradient on each worker.
-  • Gradient synchronization. The gradients of different workers are synchronized through communication primitives.
+  • Forward pass. The model parameters $x_{t}$ is applied to the input data set $X$ to obtain the prediction $Y$ as:
+    $Y=Forward(x_{t},X).$ (1)
+  • Backward pass. A loss function is used to calculate the gradient $G_{i,t}$ on $i$-th worker, satisfying that
+    $G_{i,t}=Backward(x_{t},Y).$ (2)
+    Then the gradients of workers are synchronized through communication primitives, satisfying that
+    $G_{t}=Sync(G_{i,t}).$ (3)
   • Model update.
-The optimizer (e.g., Adam [24]) uses the synchronized gradient $G_{t}$ to update the model state (denoted by $M_{t}$), which contains the model parameters $x_{t}$ and the optimizer parameters $o_{t}$ (i.e., $M_{t}=(x_{t},o_{t})$), satisfying that
-$M_{t+1}=M_{t}+Adam(G_{t}).$ (1)
+The optimizer (e.g., Adam [kingma2014adam]) uses the synchronized gradient $G_{t}$ to update the model state (denoted by $M_{t}$), which contains the model parameters $x_{t}$ and the optimizer parameters $o_{t}$ (i.e., $M_{t}=(x_{t},o_{t})$), satisfying that
+    $M_{t+1}\leftarrow M_{t}+Adam(G_{t}).$ (4)
 Here, we specially introduce the widely used optimizer, Adam.
 The Adam optimizer maintains the first-order and second-order moments (i.e., the mean and variance of the gradients), and each has the same size as the model parameters, denoted by $\Psi$.
-Therefore, Adam requires an additional $2\Psi$ parameter storage beyond the model itself, which will support one of our insights in §III-B.
+Therefore, Adam requires an additional $2\Psi$ parameter storage beyond the model itself, which will support one of our findings in §III-B.
 II-B Checkpointing Techniques
 When training large models, failures frequently occur in distributed training systems due to extensive computational resources and long training times.
-To handle frequent training failures, checkpointing, a widely used technique in common training frameworks (e.g., PyTorch [39] and Tensorflow [1]), has been extensively studied [15, 36, 54, 19, 53, 8, 32].
+To handle frequent training failures, checkpointing, a widely used technique in common training frameworks (e.g., PyTorch [paszke2019pytorch] and Tensorflow [abadi2016tensorflow]), has been extensively studied [eisenman2022check, mohan2021checkfreq, wang2023gemini, gupta2024just, wang2024fastpersist, chen2023cost, maurya2024datastates].
 It regularly copies the model state to local storage or remote storage during training.
 During recovery from the failure, all GPUs load the latest checkpoint and rerun the remaining process until the point of failure.
-To evaluate checkpointing performance, recent studies (e.g., [19]) define the wasted time metric as the sum of the recovery time from the latest checkpoint and the steady-state overhead (i.e., the GPU time for checkpointing when no failures have happened).
+To evaluate checkpointing performance, recent studies (e.g., [gupta2024just]) define the wasted time metric as the sum of the recovery time from the latest checkpoint and the steady-state overhead (i.e., the GPU time for checkpointing when no failures have happened).
 Frequent checkpointing.
-Traditional checkpointing is performed at the granularity of epochs, but an epoch may run for hours, thus incurring substantial wasted time for every failure [36].
-Thus, state-of-the-art studies [36, 54] that aim for fast recovery increase the checkpointing frequency from epochs to iterations, which we call frequent checkpointing in this paper.
+Traditional checkpointing is performed at the granularity of epochs, but an epoch may run for hours, thus incurring substantial wasted time for every failure [mohan2021checkfreq].
+Thus, state-of-the-art studies [mohan2021checkfreq, wang2023gemini] that aim for fast recovery increase the checkpointing frequency from epochs to iterations, which we call frequent checkpointing in this paper.
 The state-of-the-art systems are specified as follows:
-Checkfreq [36] decouples checkpointing into snapshot and persist operations, and pipelines them with computation to enable frequent checkpointing.
+Checkfreq [mohan2021checkfreq] decouples checkpointing into snapshot and persist operations, and pipelines them with computation to enable frequent checkpointing.
 Specifically, the snapshot operation copies model parameters from GPU to CPU memory, while the persist operation asynchronously writes the snapshot to persistent storage.
 This design allows Checkfreq to create a checkpoint every 14–19 iterations.
-Gemini [54] improves frequent checkpointing with the help of the CPU memory of the host machine and proposes a checkpointing traffic scheduling algorithm to mitigate the impact on model training.
+Gemini [wang2023gemini] improves frequent checkpointing with the help of the CPU memory of the host machine and proposes a checkpointing traffic scheduling algorithm to mitigate the impact on model training.
 In this way, Gemini increases the checkpoint frequency over Checkfreq by up to $8\times$, even per-iteration frequency.
-PCcheck [52] enables checkpointing as frequently as every 10 iterations by checkpointing to persistent main memory (PMEM).
+PCcheck [strati2025pccheck] enables checkpointing as frequently as every 10 iterations by checkpointing to persistent main memory (PMEM).
 PCcheck utilizes concurrent checkpoints and minimizes time per checkpoint using pipelining techniques and multiple threads.
-Just-in-time checkpointing [19] diverges from prior frequent checkpointing techniques.
+Just-in-time checkpointing [gupta2024just] diverges from prior frequent checkpointing techniques.
 Instead of saving model states at regular intervals, Just-in-time checkpointing initiates a checkpoint only when a failure is detected and assumes redundancy across workers to ensure fault tolerance.
@@
 Differential checkpointing.
-A recent study, Check-N-Run [15], on reducing the checkpointing costs introduces differential checkpointing (DC for short), which records the part of the model that is modified only during the last checkpoint interval so as to reduce the checkpoint size.
+A recent study, Check-N-Run [eisenman2022check], on reducing the checkpointing costs introduces differential checkpointing (DC for short), which records the part of the model that is modified only during the last checkpoint interval so as to reduce the checkpoint size.
 Specifically, DC saves its model state as full checkpoint ($C^{F}$) regularly.
-DC also saves its differential checkpoint of the $t$-th iteration (denoted by $C^{D}_{t}$) at an iteration-level high frequency.
-Then the model state $M_{n}$ at the $n$-th iteration can be reconstructed by merging the full and all differential checkpoints as
-$M_{n}=C^{F}+C^{D}_{1}+C^{D}_{2}+\dots+C^{D}_{n}.$ (2)
-However, we note that Check-N-Run implements DC only in recommendation systems, since the recommendation model has a unique property that its embedding table has a high sparsity, such that DC is particularly well suited for recommendation models where only a small part of the model parameters are updated after each iteration.
+DC also saves its differential checkpoint of the $t$-th iteration (denoted by $C^{D}_{t}$) at an iteration-level high frequency as
+$C^{D}_{t}=M_{t+1}-M_{t}.$ (5)
+Then the model state $M_{t}$ at the $t$-th iteration can be reconstructed by merging the full and all differential checkpoints as
+$M_{t}=C^{F}+C^{D}_{1}+C^{D}_{2}+\dots+C^{D}_{t-1}.$ (6)
+However, we note that Check-N-Run implements DC only in recommendation systems, since the recommendation model has a unique property that its embedding table has a high sparsity such that DC is particularly well suited for recommendation models where only a small part of the model parameters are updated after each iteration.
 In contrast, general DNN models have to be updated entirely after each iteration since gradients are computed for all the model parameters, so in this case, deploying DC has to highly compress the differentials, leading to substantial computational cost, while higher frequency incurs more cost (see §III-A).
@@
 II-C Gradient Compression Techniques
-Gradient compression techniques have been widely used in high-performance distributed training systems [3, 16, 17, 30, 31, 47, 48, 51, 13, 4, 9, 50] to significantly reduce the communication overhead of the gradient synchronization (see §II-A).
-Two representative compression approaches are Sparsification [3, 51], which selects only a subset of the elements of the gradient, and Quantization [13, 5], which reduces the number of bits of each element of the gradient.
-In this paper, we will observe that the costs of deploying DC in distributed training can be largely reduced by leveraging gradient compression (sparsification and quantization), which motivates our main idea (see §III-C for details).
+Gradient compression techniques have been widely used in high-performance distributed training systems [aji2017sparse, faghri2020adaptive, fang2019redsync, lin2017deep, m2021efficient, shi2019layer, shi2019distributed, stich2018sparsified, dettmers20158, alistarh2017qsgd, chen2021cswap, song2023optimus] to significantly reduce the communication overhead of the gradient synchronization in backward pass (see §II-A and Equation (3)).
+Two representative compression approaches are Sparsification [aji2017sparse, stich2018sparsified], which selects only a subset of the elements of the gradient, and Quantization [dettmers20158, NIPS2017QSGD], which reduces the number of bits of each element of the gradient.
+In this paper, we will observe that the costs of deploying DC in distributed training can be largely reduced by leveraging gradient compression, which motivates our main idea (see §III-C for details).
 III Observations and Motivation
@@
 As stated in §II-B, DC may incur high costs when deployed in general distributed training systems without the help of the sparsity feature in recommendation systems, so we conduct measurements to observe that show the challenges of DC in general distributed training systems, especially under high frequency.
-We conduct our measurement on eight NVIDIA A100 GPUs with a 25Gbps network [54].
-We evaluate a typical large-model training task GPT2-L [41], with a common differential checkpointing scheme (see Equation (2)).
-Detailed setups are shown in §VII-A.
+We conduct our measurement on eight NVIDIA A100 GPUs with a 25Gbps network [wang2023gemini].
+We evaluate a typical large-model training task GPT2-L [radford2019language], with a common differential checkpointing scheme (see Equation (6)).
+Detailed setups are shown in §VIII-A.
 Challenge 1: Frequent DC incurs significant computation cost for highly compressing differentials.
-In the absence of the sparsity feature’s support, DC has to apply the highly compressed scheme (compression ratio $\rho$ = 0.01 [15]) to the differentials (called compressed differentials), thus degrading the training performance.
+In the absence of the sparsity feature’s support, DC has to apply the highly compressed scheme (compression ratio $\rho$ = 0.01 [eisenman2022check]) to the differentials (called compressed differentials), thus degrading the training performance.
 To measure the impact of compression computation cost on training, Figure 1(a) shows the training time of cases without compression and with different compression frequencies (e.g., 8 iteration compression frequency means the compression is performed every 8 iterations).
-We see that compared to the case without compression, the case with compression slows down the training process by $13\%\sim 57\%$, while higher frequency incurs slower training.
+We see that compared to the case without compression, the case with compression slows down the training process greatly by $13\%\sim 57\%$, while higher frequency incurs much slower training.
 Challenge 2: Frequent DC causes significant transmission cost due to frequent checkpoint writes.
-Frequent DC needs to write the differential of model state to storage frequently, which blocks the training process frequently and incurs substantial transmission overhead, since the model update has to wait for the snapshotting of checkpoint [36, 53].
+Frequent DC needs to write the differential of model state to storage frequently, which blocks the training process frequently and incurs substantial transmission overhead, since the model update has to wait for the snapshotting of checkpoint [mohan2021checkfreq, wang2024fastpersist].
 To measure the impact of transmission cost on training in terms of frequency, Figure 1(b) shows the training time of cases without differential transmission and with different differential transmission frequencies.
@@
 Gradients of a general DNN model are used to adjust the model state via the optimizer (e.g., Adam), so based on Equation (4), we have $Adam(G_{t})=M_{t+1}-M_{t}$.
-Meanwhile, the differential checkpoint of the $t$-th iteration $C_{t}^{D}$ also equals $M_{t+1}-M_{t}$ by definition (see §II-B).
-Thus, we have $C_{t}^{D}=Adam(G_{t})$.
-Therefore, the compressed version of differential checkpoint $C_{t}^{D}$ can be directly constructed via the optimizer by the compressed gradient $\widetilde{G}_{t}$, which can be obtained in advance by the gradient compression operation.
+Meanwhile, the differential checkpoint of the $t$-th iteration $C_{t}^{D}=M_{t+1}-M_{t}$ based on Equation (5).
+Thus, we have:
+$C_{t}^{D}=Adam(G_{t}).$ (7)
+Therefore, the compressed version of the differential checkpoint $C_{t}^{D}$ can be directly constructed via the optimizer by the compressed gradient $\widetilde{G}_{t}$, which can be obtained in advance by the gradient compression operation.
 Finding 2: The compressed gradients are one-third of the compressed differentials in size.
@@
 Figure 2(a) illustrates that during a 3-iteration training, DC has three compression stalls and three transmission stalls, hindering the training process.
-Findings 1 and 2 show that the compressed gradient can construct the compressed differentials and its size is much smaller than the compressed differential.
-Thus, the compressed gradient can be leveraged to act as the compressed differential checkpoint of DC, so as to 1) remove the compression stalls via eliminating DC’s differential compression, and 2) reduce or even remove the transmission stalls since the small-size compressed gradient is likely to enable DC’s transmission to be overlapped with the training process.
+Findings 1 and 2 show that the compressed gradient can construct the compressed differentials, and its size is much smaller than the compressed differential.
+Thus, the compressed gradient can be leveraged to act as the compressed differential checkpoint of DC, so as to (1) remove the compression stalls via eliminating DC’s differential compression, and (2) reduce or even remove the transmission stalls since the small-size compressed gradient is likely to enable DC’s transmission to be overlapped with the training process.
 Therefore, we propose LowDiff, a low-cost differential checkpointing framework, whose main idea is to reuse the compressed gradients to act as the differential checkpoints with no DC compression cost and with a lower DC transmission cost.
 Figure 2(b) illustrates that during a 3-iteration training, LowDiff has no compression and transmission stalls by reusing $\widetilde{G}_{t}$, $\widetilde{G}_{t+1}$ and $\widetilde{G}_{t+2}$ as three differential checkpoints, speeding up the training.
-III-D Data dependency analysis for LowDiff
-Requirement of LowDiff.
-LowDiff requires that gradient reuse be decoupled from training, allowing DC to run in parallel.
-Without this decoupling, gradient reuse may introduce additional stalls in training (i.e., data dependency [53]).
-Specifically, in standard training, checkpointing can run independently of the forward and backward passes.
-However, checkpointing is data dependent on the model update pass [53], since it must read the parameter after model updates (i.e., Write-after-Read (WAR) [21]), which incurs a training stall for synchronization.
-Prior work [36, 15] also takes the data dependencies into account to design a checkpointing strategy that overlaps with training and reduces stalls.
-Consequently, it is crucial to carefully examine the coupling introduced by gradients between training and checkpointing.
-Next, we first show the data dependency when deploying DC directly, and then explore how reusing compressed gradients for DC interacts with the training process.
-Deploying DC directly introduces additional data dependencies.
-As shown in Figure 3(a), checkpointing depends not only on the latest model state but also on the model state from $n$ iterations earlier.
-This introduces two major issues: 1) The model update in iteration $t+1$ must wait for checkpointing in iteration $t$, leading to a training stall due to the transmission of checkpointing.
-2) The model state from iteration tnt-n must be retained in GPU memory until the differential computation in iteration $t$, resulting in extra GPU memory overhead.
-Moreover, both checkpoint compression and writing lie on the critical path of these data dependencies, and therefore cannot be effectively overlapped with training.
-Reusing compressed gradients for DC eliminates such data dependency.
-Gradients are generated during the backward pass and remain unchanged thereafter.
-The update pass only reads these gradients (read-only) to update the model state, which removes the Write-after-Read (WAR) dependency and allows gradient reuse to be performed immediately after backpropagation without waiting for the model update.
-Moreover, by treating the compressed gradient as the compressed differential, the construction of the differential no longer depends on the model states from iterations $t$ or tnt-n, eliminating prior data dependencies for DC calculation.
-As illustrated in Figure 3(b), together these factors ensure that gradient reuse introduces no data dependency from checkpointing to training, thereby decoupling the two processes and allowing them to execute fully in parallel.
-Figure 3: Data dependency analysis.
-IV Design
+While LowDiff provides a low-cost DC way based on gradient reuse, its integration into training systems, particularly the feasibility of concurrent execution with the training process, requires further analysis.
+We next analyze such concurrency in §IV.
+IV Analysis
+By incorporating the main idea of gradient reuse (§III), LowDiff aims to enable the DC process to run concurrently with model training, thereby minimizing potential stalls.
+To confirm the feasibility of such parallelism, we analyze our proposed strategy from two key aspects:
+  • Parallelism property: it ensures no data dependencies block concurrency with training iterations.
+  • Execution time overlap: it reveals whether the checkpointing cost can be effectively hidden within the training time.
+We provide a detailed analysis of these two aspects as follows.
+Figure 3: LowDiff’s DC can run in parallel with forward pass(F), backward pass(B), and model update(U).
+IV-A Parallelism Analysis
+Necessity of Parallelism Analysis.
+Checkpointing can be executed with the forward and backward passes which do not modify the model state [mohan2021checkfreq], but it cannot run in parallel with model update which directly writes to it [wang2024fastpersist].
+In other words, checkpointing is data-dependent on the model update [wang2024fastpersist].
+The reason is that it must read the model state after model updates (Read-after-Write [hennessy2011computer], RAW dependency), and the model update at the next iteration must write the model state after checkpointing completes its read (Write-after-Read [hennessy2011computer], WAR dependency).
+Consequently, the two dependency patterns incur training stalls, preventing from paralleling checkpointing with model update.
+Prior studies [mohan2021checkfreq, wang2023gemini, wang2024fastpersist] design their checkpointing strategies based on these data dependencies to pipeline training and checkpointing.
+Consequently, it is necessary to carefully analyze: (1) the data dependency introduced by our reuse strategy and (2) the parallelism property of LowDiff’s DC, specified as follows.
+(1) LowDiff’s DC can eliminate the WAR dependency. We next analyze current DC’s WAR dependency with next iteration’s model update, backward, and forward pass.
+  • Model update: as described in §II-A, the model update (Equation (4)) only requires read-only access to the gradient.
+Because LowDiff checkpoints only the compressed gradient, leveraging gradient reuse, the original Write-after-Read (WAR) dependency is eliminated because the checkpointing process no longer needs to read the model state ($M_{t}$).
+Instead, it just reads the compressed gradient ($\widetilde{G}_{t}$), which is available in the backward pass and will not be changed after.
+Since both the checkpointing and model update are read-only consumers of the gradient, they can operate in parallel without conflict.
+  • Backward pass: LowDiff’s gradient reuse strategy introduces no additional WAR dependency on backward pass, since each gradient ${G}_{t}$ is finalized within its iteration and never modified thereafter.
+Specifically, once a gradient ${G}_{t}$ is generated in the current iteration ($t$), it is final and will not be modified by subsequent steps.
+The backward pass of the $t+1$-th iteration (Equation (2)) does not read or depend on ${G}_{t}$; it generates a completely new gradient ($G_{t+1}$) from scratch and won’t change the ${G}_{t}$.
+Therefore, the DC’s read of ${G}_{t}$ does not conflict with the $B$’s write of ${G}_{t+1}$ at $t+1$-th iteration, allowing LowDiff’s DC to run in parallel with the next iteration’s backward pass.
+  • Forward pass: The forward pass (Equation (1)) doesn’t involve gradient, so there is no WAR dependency between LowDiff’s DC and forward pass.
+(2) LowDiff can run in parallel with forward pass, backward pass, and model update.
+To visualize the parallelism benefit from data dependency, Figure 3 shows the execution timeline with LowDiff.
+We see that the DC task ($C^{D}_{t}$) at the $t$-th iteration is no longer coupled with the model update ($U$) or the following iteration (no dependency arrow from DC to training).
+Instead, the reusing operation on compressed gradient ($\widetilde{G}_{t}$) begins immediately after the backward pass ($B$), running in parallel with the model update and subsequent iteration’s forward pass(F), backward pass (B), and model update(U).
+This parallelism property effectively transforms LowDiff’s DC from a blocking operation to a parallel execution with training, thereby eliminating the synchronization stalls previously caused by the WAR dependency.
+Figure 4: Time of iteration, full checkpointing, and differential checkpointing.
+IV-B Overlap Analysis
+Necessity of Overlap Analysis.
+Beyond its parallelism property, LowDiff can perform even better if the checkpointing process is fast enough to be hidden by the training iteration, thus minimizing DC’s cost.
+Specifically, the time required to checkpoint a compressed gradient (denoted as $C$) should be shorter than a single training iteration (denoted as $F+B+U$), similar to the Microsoft study [mohan2021checkfreq].
+Consequently, it is also necessary to analyze the potential of overlapping, so we first (1) take a measurement analysis of training time, full checkpointing, and DC time, as well as (2) discuss why the overlap can be practically feasible.
+(1) The checkpointing of compressed gradients can be overlapped with training.
+Following the measurement setup in §III-A, we measure the iteration time, full checkpointing time and differential checkpointing time on BERT-B, BERT-L, GPT2-S, and GPT2-L.
+As shown in Figure 4, the DC time is only 20.5%, 22.1%, 23.1%, and 24.6% of the corresponding training iteration time for these models.
+This indicates that in most cases, the checkpointing process can be fully overlapped with training iterations without introducing noticeable stalls.
+(2) Modern hardware makes the overlap practically feasible.
+NVMe SSDs based on PCIe 4.0 offer write speeds around 5 GB/s (and up to 12 GB/s for PCIe 5.0, with RAID 0 configurations being even faster).
+Given that LowDiff’s compressed-gradient checkpoints are relatively small (several hundred MBs to a few GBs, as evaluated in §VIII), the theoretical I/O write time is typically under 1 second.
+This duration is significantly shorter than that of a full training iteration [wang2023gemini].
+Based on the above analysis, we confirm that the concurrent execution between LowDiff’s DC and training is theoretically feasible.
+While this analysis establishes the feasibility of frequent DC, the detailed design of LowDiff will be presented in §V.
+V Design
 Figure 5: Architecture of LowDiff
 We design LowDiff based on the main idea in §III-C with the following design goals.
-  • Lowering the costs of DC. LowDiff reduces or even eliminates the computation and transmission costs of compressed differential checkpoints by reusing compressed gradients (§IV-A).
-  • Enhancing DC performance. LowDiff accelerates the differential checkpointing under high frequency via batched gradient writing optimization (§IV-B).
-  • Minimizing the wasted time. LowDiff minimizes the wasted time by analytically tuning the full checkpointing frequency and gradient batching size (§IV-C).
+  • Lowering the costs of DC. LowDiff reduces or even eliminates the computation and transmission costs of compressed differential checkpoints by reusing compressed gradients (§V-A).
+  • Enhancing DC performance. LowDiff accelerates the differential checkpointing under high frequency via batched gradient writing optimization (§V-B).
+  • Minimizing the wasted time. LowDiff minimizes the wasted time by analytically tuning the full checkpointing frequency and gradient batching size (§V-C).
 Figure 5 shows the architecture of LowDiff.
@@
 Finally, the batched differential and full checkpoints are persisted to local or remote storage.
-IV-A Reusing Compressed Gradients
+V-A Reusing Compressed Gradients
 Design requirements.
@@
 In PyTorch, this can be achieved with torch.multiprocessing.Queue(), where the inter-process communication creates an IPC handle when a tensor is enqueued and reopens it when dequeued.
+Algorithm 1 Reusing in LowDiff
+1: Model state $M_{t}$ at iteration $t$, Gradient $G_{i,t}$ at worker $i$, Reusing queue $Q$, Full checkpoint $C^{F}_{t}$, Differential checkpoint $C^{D}_{n}$ at iteration $n$, Compressor CompComp, Decompressor $Comp^{-1}$.
+2: function Training Process
+3:   for all $t$ do
+4:    $G_{i,t}\leftarrow Backward(x_{t})$
+5:    $\widetilde{G}_{i,t}\leftarrow Comp(G_{i,t})$ $\triangleright$ Compress
+6:    $\widetilde{G}_{t}\leftarrow Sync(\widetilde{G}_{i,t})$ $\triangleright$ Synchronize
+7:    ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(\widetilde{G}_{t})$ $\triangleright$ Zero-Copy
+8:    $G_{t}\leftarrow Comp^{-1}(\widetilde{G}_{t})$ $\triangleright$ Decompress
+9:    $M_{t+1}\leftarrow M_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update
+10: function Checkpointing Process
+11:   while $\widetilde{G}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Reusing
+12:    $C^{D}_{t}\leftarrow Save(\widetilde{G}_{t})$ $\triangleright$ Diff CKPT
+13:   $C^{F}_{t}\leftarrow Save(M_{t})$ $\triangleright$ Full CKPT Regularly
+14: function Recovery Process
+15:   $M_{t}\leftarrow Load(C^{F}_{t})$ $\triangleright$ Load Full CKPT
+16:   for $j\leftarrow t$, …, $n$ do $\triangleright$ Restore to Latest
+17:    $\widetilde{G}_{j}\leftarrow Load(C^{D}_{j})$ $\triangleright$ Load Diff CKPT
+18:    $G_{j}\leftarrow Comp^{-1}(\widetilde{G}_{j})$ $\triangleright$ Decompress
+19:    $M_{j+1}\leftarrow M_{j}+Adam(G_{j})$ $\triangleright$ Diff Merge
+20:   return $M_{n}$
 Design details Algorithm 1 specifies the following details of reusing compressed gradients, including training, checkpointing, and recovery processes.
@@
 The checkpointing process repeatedly retrieves the compressed gradients from Reusing Queue $Q$ to CPU memory (Line 11).
-Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §IV-B).
+Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §V-B).
 After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12-14) to storage.
@@
 By repeating this process for $n$ iterations (Lines 18-23), where compressed gradients from differential checkpoints are successively loaded and applied, the model state is finally restored to $M_{n}$ (Line 24), thus recovering the training from failures.
-Algorithm 1 Reusing in LowDiff
-Model state $M_{t}$at iteration $t$, Gradient $G_{i,t}$at worker $i$, Reusing queue $Q$, Full checkpoint $C^{F}_{t}$, Differential checkpoint $C^{D}_{n}$at iteration $n$, Compressor CompComp, Decompressor $Comp^{-1}$.
-Training Process $t$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-IV-B Batched Gradient Writing Optimization
-Limitations of LowDiff.
-We find that LowDiff in §IV-A still has two limitations: Limitation 1: Many writes due to frequent DC.
-The high frequency leads to numerous differential checkpoints, which incur significant write overhead.
-Limitation 2: Increased GPU memory consumption from frequent gradient reuse.
-The high frequency of gradient reuse results in many compressed gradients being retained in GPU memory, waiting to be written to storage.
+V-B Batched Gradient Writing Optimization
+Limitations of LowDiff. We find that LowDiff in §V-A still has two limitations:
+  • Limitation 1: Many writes due to frequent DC. The high frequency leads to numerous differential checkpoints, which incur significant write overhead.
+  • Limitation 2: As analyzed in §IV-A, avoiding the WAR dependency requires temporarily retaining the compressed gradient $\widetilde{G}_{t}$ for the DC operation.
+Consequently, frequent DCs cause this gradient buffer to remain occupied while waiting for the write to storage, leading to extra GPU memory consumption.
+we provide the batched write and offloading strategy to address the corresponding limitation as follows.
 Batched write.
-To address Limitation 1, we can batch multiple writes of compressed gradients (reused as differential checkpoints) into a single write operation, with the help of the widely-used gradient accumulation technique [30, 22, 2] where the gradients of the same shape and size can be accumulated.
+To address Limitation 1, we can batch multiple writes of compressed gradients (reused as differential checkpoints) into a single write operation, with the help of the widely-used gradient accumulation technique [lin2017deep, hermans17a, accumulation] where the gradients of the same shape and size can be accumulated.
 Note that the above batching scheme, which benefits DC writes, cannot work for traditional checkpointing that only saves full checkpoints, since if a new full checkpoint comes, the previous full checkpoints become obsolete and can be discarded without the need to accumulate them.
 Offloading batching to CPU.
-To address Limitation 2, we adopt a classic GPU-CPU offloading scheme [42] that offloads the batching operation to the CPU and caches the compressed gradients in CPU memory, thus reducing the GPU memory cost caused by frequent checkpointing.
+To address Limitation 2, we adopt a classic GPU-CPU offloading scheme [rajbhandari2020zero] that offloads the batching operation to the CPU and caches the compressed gradients in CPU memory, thus reducing the GPU memory cost caused by frequent checkpointing.
 Note that the batching operation can be easily handled by the CPU since the operation mainly involves the addition of compressed gradients.
-Design details To solve Limitation 1 & 2, we design a batched gradient writing optimization, which offloads multiple compressed gradients from GPU to CPU memory, and then aggregates them into batched writes to storage.
+Figure 6: Batched gradient write optimization
+Design details.
+To solve Limitation 1 & 2, we design a batched gradient writing optimization, which offloads multiple compressed gradients from GPU to CPU memory, and then aggregates them into batched writes to storage.
 As shown in Figure 6, we specify the batched gradient writing optimization of LowDiff in three steps:
@@
 The checkpointing process extracts the batched differential checkpoint $C^{B}_{t}$ from the CPU buffer and writes it to storage in a single I/O operation.
-Figure 5: Batched gradient write optimization
-IV-C Optimizing Checkpointing Configuration
+V-C Optimizing Checkpointing Configuration
 Insight into configuring LowDiff.
-Note that recent studies on optimizing the checkpointing performance only focus on configuring the checkpointing frequency (e.g., [19]).
-However, we observe an insight that LowDiff’s checkpointing performance will also rely on the batching size, i.e., the amount of the batched gradients every batching, introduced in §IV-B.
+Note that recent studies on optimizing the checkpointing performance only focus on configuring the checkpointing frequency (e.g., [gupta2024just]).
+However, we observe an insight that LowDiff’s checkpointing performance will also rely on the batching size, i.e., the amount of the batched gradients every batching, introduced in §V-B.
 To verify the insight, we take another measurement analysis with the same setting as that in §III-A.
@@
 Note that all the results are normalized by comparing them to the lowest wasted time.
-We first see that when FCF is too low ($FCF=10$) or too high ($FCF=100$), the wasted time will be increased, since a low FCF will incur high recovery time while a high FCF will incur high steady state overhead (GPU time), Thus, we need to tune FCF to minimize the wasted time, which has been found by Microsoft [19].
+We first see that when FCF is too low ($FCF=10$) or too high ($FCF=100$), the wasted time will be increased, since a low FCF will incur high recovery time while a high FCF will incur high steady state overhead (GPU time), Thus, we need to tune FCF to minimize the wasted time, which has been found by Microsoft [gupta2024just].
 More importantly, under the same FCF (i.e., in the same row), we also need to tune the batching size (BS for short) for minimizing the wasted time, For example, when $FCF=10$ or $20$, the wasted time is minimal when $BS=2$; when $FCF=50$ or $100$, the wasted time is minimal when $BS=3$.
@@
 We will use $(f^{*},b^{*})$ as the optimal configuration in LowDiff.
-V enhancement
+VI enhancement
+The current design of LowDiff relies on gradient compression to enable efficient gradient reuse.
+However, in some high-precision training scenarios, any potential accuracy degradation caused by compression is unacceptable.
+If the distributed training system does not inherently employ a gradient compression strategy, making it impossible to leverage compression-induced bandwidth advantages.
+So we also need to design a non-compression high-frequency DC strategy for above scenarios.
+We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
 Figure 7: Layer-wise gradient reuse.
-Issue of LowDiff without gradient compression.
-The current design of LowDiff relies on gradient compression to enable efficient gradient reuse.
-When compression is disabled, we can still enforce a lightweight, reuse-only compression to preserve the bandwidth reduction in Finding 2, but the computational savings of Finding 1 no longer apply.
-In this case, the differential checkpoints become significantly larger, slowing down their snapshotting and persistence accordingly.
-As a result, checkpointing competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall introduced by gradient reuse, snapshotting, and persistence of DC under such non-compressed scenarios.
-We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
-V-A Layer-wise gradient reusing and snapshotting.
-A key challenge in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of blocking.
+VI-A Layer-Wise Gradient Reusing and Snapshotting Strategy.
+Issue of snapshotting.
+A key issue in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of a stall.
+The snapshotting competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall.
 To accelerate such large-scale snapshotting, we aim to design a fine-grained strategy that better overlaps gradient transmission with training.
-As shown in Figure 6, gradients are produced incrementally in the backward pass, following the reverse layer order during training.
-Instead of waiting until the entire backward pass finishes to snapshot all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [10], PyTorch DDP [28], and Horovod [46]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced.
-Inspired by this execution model, we design a fine-grained reuse strategy: immediately snapshot each layer’s gradient to host memory upon computation and stream it into a differential checkpoint assembler.
-By pipelining backpropagation, communication, and gradient reuse at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
+Insight 1: gradient can be reused layer-by-layer.
+We recall that gradients are produced incrementally in the backward pass, following the reverse layer order during training, as shown in Figure 7.
+Instead of waiting until the entire backward pass finishes to synchronize all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [deepspeed], PyTorch DDP [li2020pytorch], and Horovod [sergeev2018horovod]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced.
+Inspired by this execution model, we design a fine-grained reuse strategy: once a layer’s gradient has been synchronized, we immediately reuse it to the reusing queue, snapshot it to host memory, and stream it into a differential checkpoint assembler.
+This granular approach enables parallel data movement, allowing concurrent snapshot writes to CPU DRAM to better utilize memory bandwidth, thereby mitigating the high transfer volume.
+By pipelining backpropagation, communication, gradient reuse, and snapshotting at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
 Figure 8: Batching strategies of LowDiff and LowDiff+.
-V-B CPU-based asynchronous persistence strategy.
-The layer-wise strategy addresses the snapshotting issue in LowDiff with compressed gradients.
-Another challenge lies in the persistence of both full and differential checkpoints.
-The batching strategy described in §IV-B accumulates small compressed updates on the CPU side before writing them out.
+VI-B CPU-Based Asynchronous Persistence Strategy.
+Issue of persistence.
+While the layer-wise strategy addresses the snapshotting issue in the non-compression setting, another issue remains: persistence of both full and differential checkpoints.
+The batching strategy described in §V-B accumulates small compressed updates on the CPU side before writing them out.
 It also periodically persists the full model state, which works effectively when update sizes are small and sparse.
 In the non-compression setting, differential checkpoints are significantly larger, and naively batching them demands a sustained write throughput that exceeds the available storage bandwidth, making it essential to design a more efficient checkpoint merging strategy that better lowers the writing data volume.
-To address this, we observe that in the non-compression setting, gradients do not need to be persisted separately.
+Insight 2: differential and full checkpoints can be batched together in CPU memory.
+We find that in the non-compression setting, the differential checkpoints (i.e.
+gradients) and full checkpoints do not need to be persisted separately.
 Instead, they can be directly fused with the full model state inside CPU memory, enabling the CPU to maintain an always-up-to-date replica of the GPU model.
@@
 By fusing gradients into the CPU-resident model state instead of persisting them separately, redundant writes of differential checkpoints are avoided, which directly reduces the overall storage volume.
-The different batching strategies of LowDiff and LowDiff+ are shown in 7.
+The different batching strategies of LowDiff and LowDiff+ are shown in Figure 8.
 This strategy decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
-V-C Design of LowDiff+
+VI-C Design of LowDiff+
 Design details.
-Based on Insight 1 and 2, we design LowDiff+ without gradient compression.
+Based on Insight 1 and 2, we design LowDiff+ under non-compression scenarios.
 The architecture of LowDiff+ is shown in Figure 9, and its detailed algorithm is presented in Algorithm 2.
@@
 Once the communication in a thread finishes, the corresponding gradients are forwarded via the reusing queue to the CPU-side checkpointing process (Line 16), while the training process waits for the full gradient $G_{t}$ (i.e., wait for handle set $H_{g}$) (Line 6) to update the GPU model $M^{G}_{t}$ (Line 7).
-Figure 8: Workflow of LowDiff without gradient compression.
 In the checkpointing process, gradients $G^{n}_{t}$ streamed from the GPU are reused layer by layer (Line 9).
 The offloading thread pool $P_{s}$ handles multiple layers in parallel (Line 10), snapshotting the layer-wise gradients into CPU memory as they become available (Line 19).
+This parallel snapshotting allows different layer gradients to be written concurrently to CPU memory, thereby better utilizing the available DRAM bandwidth.
 We initialize a CPU-resident replica $M^{C}_{t}$ of the model parameters, aligned with the GPU state.
 After snapshotting (Line 11), reused gradients are applied to $M^{C}_{t}$ to keep it synchronized with the GPU model $M^{G}_{t}$ (Line 12).
-The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [54].
+While gradients are snapshotted layer-by-layer, the CPU-side update (Line 12) is performed once the full gradient is aggregated.
+This is because the Adam optimizer requires the complete gradient to correctly update its internal states (e.g., momentum and variance).
+The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [wang2023gemini].
 Since such updates do not involve forward or backward computations, their execution on the CPU introduces negligible overhead.
 Persistence is performed asynchronously by checkpointing the CPU-maintained model replica (Line 13), rather than the raw gradients.
-The frequency of persistence is decided similarly to Checkfreq [36], which ensures maximal overlapping between the persist and update in CPU.
+The frequency of persistence is decided similarly to Checkfreq [mohan2021checkfreq], which ensures maximal overlapping between the persist and update of the model in CPU.
 This design both avoids redundant storage of differential checkpoints and completely decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
-When failure occurs, prior work [54, 19] categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures.
+Figure 9: Architecture of LowDiff+.
+Recovery of LowDiff+.
+When failure occurs, prior work categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures.
 For hardware failures (e.g., GPU device crashes), the distributed training system must replace the failed machine with a new one.
 In this case, the in-memory training state is lost, so the recovery process must reload model states from persistent storage checkpoints, which can be time-consuming and increase wasted time.
-In contrast, software failures (e.g., software bugs) only cause the training process to terminate while leaving the CPU memory state of the checkpointing process intact.
+In contrast, software failures (e.g., software bugs) only cause the training process to terminate.
+This design assumes the failure is confined to the training processes (e.g., a CUDA error or NaN loss) and does not corrupt the separate checkpointing process’s memory, thus leaving the CPU memory state of the checkpointing process intact.
 In such cases, recovery is achieved by reinitializing the training process and restoring the preserved CPU-resident model state to the GPUs, leveraging in-memory checkpoints.
@@
 Algorithm 2 Reusing in LowDiff+
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-VI Implementation
-VI-A Implementation of LowDiff
-We implement LowDiff atop PyTorch [40, 39] and DeepSpeed [10, 44], consisting of three main modules: reusing queue (§IV-A), batched gradient writing (§IV-B), and optimal configuration (§IV-C).
+1: GPU model state after the $t$-th iteration $M^{G}_{t}$, gradient of the $n$-th layer on node $i$ $G^{n}_{i,t}$, gradient communication and offload thread pools $P_{g}$ and $P_{s}$, reusing queue $Q$, corresponding handle sets $H_{g}$ and $H_{s}$.
+2: function Training Process
+3:   for all $t$ do
+4:    for all $n$ do
+5:      $G^{n}_{i,t}\leftarrow Backward(x^{n}_{t})$
+6:      $H_{g}.append(P_{g}.execute(\textsc{Sync},G^{n}_{i,t}))$ 7:                                         $\triangleright$ Sync thread
+8:    $G_{t}\leftarrow H_{g}.wait()$ $\triangleright$ Wait for sync
+9:    $M^{G}_{t+1}\leftarrow M^{G}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in GPU
+10: function Checkpointing Process
+11:   while $G^{n}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Zero-copy reuse
+12:    $H_{s}.append(P_{s}.execute(\textsc{Snapshot},G^{n}_{t}))$ 13:                                       $\triangleright$ Snapshot thread
+14:   $G_{t}\leftarrow H_{s}.wait()$ $\triangleright$ Wait for offload
+15:   $M^{C}_{t+1}\leftarrow M^{C}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in CPU
+16:   $C^{F}_{t}\leftarrow Save(M^{C}_{t})$ $\triangleright$ Persistence fully
+17: function Sync Thread
+18:   $G^{n}_{t}\leftarrow Sync(G^{n}_{i,t})$ $\triangleright$ Synchronization
+19:   ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(G^{n}_{t})$ $\triangleright$ Zero-copy reuse
+20:   return $G^{n}_{t}$
+21: function Snapshot Thread
+22:   $G^{n}_{t}\leftarrow Offload\_to\_cpu(G^{n}_{t})$ $\triangleright$ Snapshot
+23:   return $G^{n}_{t}$
+VII Implementation
+VII-A Implementation of LowDiff
+We implement LowDiff atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed], consisting of three main modules: reusing queue (§V-A), batched gradient writing (§V-B), and optimal configuration (§V-C).
 In addition, a parallel recovery module accelerates recovery from a large number of differential checkpoints at high frequency.
@@
 Starting from a default setup, it adapts to runtime metrics (failure rate, bandwidth) using stepwise adjustments to balance steady-state overhead and recovery cost, achieving efficient fault-tolerant checkpointing.
-Parallel recovery module.Traditional recovery methods for differential checkpointing [15] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (2)), whose the number of merge operations is $n$ with given $n$ differential checkpoints.
+Parallel recovery module.Traditional recovery methods for differential checkpointing [eisenman2022check] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (6)), whose the number of merge operations is $n$ with given $n$ differential checkpoints.
 Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints.
@@
 Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
-VI-B Implementation of LowDiff+
-LowDiff+ is also implemented atop PyTorch [40, 39] and DeepSpeed [10, 44].
+Figure 10: Parallel Fast Recovery
+VII-B Implementation of LowDiff+
+LowDiff+ is also implemented atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed].
 Since the checkpointing process is spawned from the training process, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning.
@@
 Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation.
-Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [36], thereby maximizing checkpointing throughput without interfering with GPU training.
-Figure 9: Parallel Fast Recovery
-VII Evaluation
+Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [mohan2021checkfreq], thereby maximizing checkpointing throughput without interfering with GPU training.
+VIII Evaluation
 In this section, we evaluate the efficiency of LowDiff to answer the following questions:
@@
 TABLE II: Experimental setup
-VII-A Experimental Setup
+VIII-A Experimental Setup
 Servers.
@@
 Workloads.
-We evaluate widely-used DNN models on tasks of image classification and natural language processing (NLP), including ResNet [20], VGG [49], Bert [14], and GPT-2 [41].
-The image classification models are evaluated on the CIFAR-100 [25] and ImageNet [12] datasets, and NLP models are evaluated on the SQuAD [43] and WikiText [33] datasets.
+We evaluate widely-used DNN models on tasks of image classification and natural language processing (NLP), including ResNet [he2016deep], VGG [simonyan2014very], Bert [devlin2018bert], and GPT-2 [radford2019language].
+The image classification models are evaluated on the CIFAR-100 [krizhevsky2009learning] and ImageNet [deng2009imagenet] datasets, and NLP models are evaluated on the SQuAD [SQuAD] and WikiText [merity2016pointer] datasets.
 The detailed configurations are shown in Table II(b).
-We use Adam [24] as the default optimizer.
-The training process uses sparsification compression with a common compress ratio $\rho$ of 0.01 [3, 29, 18, 35, 7, 45].
+We use Adam [kingma2014adam] as the default optimizer.
+The training process uses sparsification compression with a common compress ratio $\rho$ of 0.01 [aji2017sparse, li2022near, fei2021efficient, ming2024adtopk, chen2023bbtopk, renggli2019sparcml].
 We use default batching sizes and learning rates according to the original papers or standard practices.
-We also implement the LowDiff system on the VGG-16 model with pipeline parallelism provided in the DeepSpeedExamples [11].
+We also implement the LowDiff system on the VGG-16 model with pipeline parallelism provided in the DeepSpeedExamples [deepspeedexample].
 Baselines.
-We evaluate two state-of-the-art checkpointing strategies: CheckFreq [36] and Gemini [54], as described in §II-B.
+We evaluate two state-of-the-art checkpointing strategies: CheckFreq [mohan2021checkfreq] and Gemini [wang2023gemini], as described in §II-B.
 The experiments follow the default configurations according to their original papers.
-Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [15] does in recommendation systems.
-VII-B Performance Evaluation
+Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [eisenman2022check] does in recommendation systems.
+VIII-B Performance Evaluation
 Figure 11: (Exp. 1) Training time of different models with different checkpointing strategies.
@@
 Figure 12 shows that the training time of LowDiff+ across all tasks exhibits a slight increase compared to the W/O CKPT baseline, yet it remains significantly lower than that of other checkpointing solutions.
-Specifically, LowDiff+’s training time increases by 8.2%–10.1% relative to W/O CKPT, which is slightly higher than the increase observed in the setting without gradient compression.
-This overhead is primarily due to the frequent and large-volume gradient transfers between GPU and CPU memory, which occupy PCIe bandwidth and consequently reduce training performance.
-In the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods.
+Specifically, LowDiff+’s training time increases by 7.2%–9.1% relative to W/O CKPT.
+As expected, this overhead is higher than the 2.4%-3.1% observed for LowDiff in Exp.
+1.
+This increase represents the necessary and expected performance trade-off for handling uncompressed, high-precision gradients, which was the primary motivation for the LowDiff+ design.
+The overhead is mainly due to frequent and large-volume gradient transfers between GPU and CPU memory, occupying PCIe bandwidth.
+Despite this minor overhead, in the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods.
 For the GPT2-L training task, it reduces the training time by 51.8% compared to Gemini and by 81.7% compared to CheckFreq, demonstrating its superior efficiency in settings where frequent non-compression checkpoints are required.
-Figure 12: (Exp. 3) Wasted time under different MTBFs.
 Experiment 3 (Wasted time).
@@
 Failures were simulated during training by adhering to a fixed Mean Time Between Failures (MTBF) metric.
-Using three MTBF values at 0.5, 1, and 2 [19], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S.
-Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §IV-C.
-For LowDiff+, we evaluate its wasted time under both software and hardware failures (LowDiff+ (S) and LowDiff+ (H), respectively).
+Using three MTBF values at 0.5, 1, and 2 [gupta2024just], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S.
+Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §V-C.
+For LowDiff+, we evaluate its wasted time under both software and hardware failures, where the corresponding checkpoints can be accessed throught snapshot and persist ( LowDiff+ (S) and LowDiff+ (P), respectively).
 Figure 13 shows that the wasted time gap between LowDiff and other checkpointing methods keeps increasing as failures become more frequent, while LowDiff maintains the lowest wasted time.
@@
 6.
-Figure 13: (Exp. 4) Maximum checkpointing frequency.
+Figure 13: (Exp. 3) Wasted time under different MTBFs.
 Experiment 4 (Checkpointing frequency).
-To evaluate the maximum checkpointing frequency that each method can achieve under bounded training speed (set as 3.5%, as the degraded speed of model training is required to be less than 3.5% in Microsoft’s study [36]), we evaluate the highest checkpointing frequency of different methods.
+To evaluate the maximum checkpointing frequency that each method can achieve under bounded training speed (set as 3.5%, as the degraded speed of model training is required to be less than 3.5% in Microsoft’s study [mohan2021checkfreq]), we evaluate the highest checkpointing frequency of different methods.
 We denote LowDiff+ ’s in-memory checkpointing as LowDiff+ (S) and its persistence as LowDiff+ (P).
@@
 These results validate our motivation that LowDiff can support high-frequency checkpointing without stalling training progress.
+Figure 14: (Exp. 4) Maximum checkpointing frequency.
 Figure 15: (Exp. 5) Recovery time of different methods.
@@
 We evaluate the recovery time of different recovery strategies under different full checkpointing frequency in GPT2-S.
-We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [39]), Naïve DC, LowDiff with parallel recovery (see §VI), and LowDiff+ from software failures (LowDiff+ (S) for short).
+We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [paszke2019pytorch]), Naïve DC, LowDiff with parallel recovery (see §VII), and LowDiff+ from software failures (LowDiff+ (S) for short).
 Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
@@
 Experiment 6 (Checkpointing time reduction and GPU memory reduction).
-To evaluate the reduction in average checkpointing time achieved by the batched write optimization proposed in §IV-B, we conduct experiments on various models with different batching sizes.
+To evaluate the reduction in average checkpointing time achieved by the batched write optimization proposed in §V-B, we conduct experiments on various models with different batching sizes.
 Figure 16a shows that the batched write strategy significantly reduces the average time required to write differential checkpoints, with a maximum reduction of $30.9\%$ when the batching size is 20 during GPT2-S training.
 The batching strategy effectively accelerates the checkpoint writing process.
-After tuning the batching frequency with the configuration strategies discussed in §IV-A, the optimization can adapt to different compression ratios and minimize the impact on training performance.
-Additionally, we also evaluate the reduction of GPU memory achieved by the batching offloaded in CPU §IV-B.
+After tuning the batching frequency with the configuration strategies discussed in §V-A, the optimization can adapt to different compression ratios and minimize the impact on training performance.
+Additionally, we also evaluate the reduction of GPU memory achieved by the batching offloaded in CPU §V-B.
 Figure 16b shows that the GPU memory usage will increase by $10\%-12\%$ without offloaded batching, especially for GPT2-L, potentially leading to memory exhaustion.
@@
 Compared to Naïve DC, LowDiff further reduces the storage overhead by 90.5%.
-Because LowDiff utilizes compressed gradients to construct differentials for both model parameters and optimizer parameters (Finding 2 in §III-B), while Naïve DC (Check-N-Run [15]) does not apply sparsification compression to the optimizer parameters within differential checkpoints, resulting in the optimizer occupying the majority of the checkpoint space.
+Because LowDiff utilizes compressed gradients to construct differentials for both model parameters and optimizer parameters (Finding 2 in §III-B), while Naïve DC (Check-N-Run [eisenman2022check]) does not apply sparsification compression to the optimizer parameters within differential checkpoints, resulting in the optimizer occupying the majority of the checkpoint space.
 The result ensures our motivation that LowDiff can significantly reduce the checkpoint size and thus lower the storage cost.
@@
 We evaluate the impact of different compression ratios $\rho$ on the checkpoint frequency of GPT2-L and GPT2-S training.
-The compression ratio $\rho$ varies from 0.1 to 0.001, consistent with the ranges commonly used in prior work [29, 35].
+The compression ratio $\rho$ varies from 0.1 to 0.001, consistent with the ranges commonly used in prior work [li2022near, ming2024adtopk].
 As shown in Figure 17, LowDiff enables per-iteration checkpointing for GPT2-S training across the entire range of $\rho\in[0.001,0.1]$ due to its smaller enough gradient size for checkpointing to be overlapped with training within one iteration.
@@
 The result shows LowDiff ensures frequent checkpointing (intervals ¡ 3 iterations) under the common compression ratio.
-VII-C System Scalability
+VIII-C System Scalability
 We test the scalability of LowDiff and LowDiff+ on V100 servers in scenarios with frequent failures and more GPUs to demonstrate the robustness and efficiency of LowDiff and LowDiff+.
@@
 We conduct experiments on the performance of LowDiff and LowDiff+ in extreme scenarios characterized by more frequent failures of the training process with MTBF ranging from 0.1 to 5 hours.
-Using the effective training time ratio metric as defined in Gemini [54], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
-Figure 17 demonstrates that LowDiff consistently sustains a higher effective training time ratio across all tested scenarios, especially under environments with frequent failures.
-Although LowDiff+ achieves slightly lower ratios than LowDiff, it still outperforms the other methods.
-For example, when the MTBF is set to 0.3 hours, LowDiff reaches an effective training time ratio of 92%, while LowDiff+ achieves 86%, compared to only 81% for Gemini and 76% for Checkfreq.
-These results highlight the robustness and efficiency of both LowDiff and LowDiff+ in handling frequent failures during training.
+Using the effective training time ratio metric as defined in Gemini [wang2023gemini], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
+Note that LowDiff+ has two recovery modes: LowDiff+ (S) for software failures (recovering from CPU memory) and LowDiff+ (P) for hardware failures (recovering from persistent storage).
+Figure 18 demonstrates that LowDiff+ (S) consistently achieves the highest effective training time ratio across all scenarios, especially under environments with frequent failures.
+This is because its near-instantaneous recovery from in-memory checkpoints (as shown in Exp.
+5) almost completely eliminates recovery overhead, compensating for its slightly higher steady-state overhead compared to LowDiff.
+LowDiff achieves the second-highest ratio.
+LowDiff+ (P) also consistently outperforms the other baselines, though its reliance on slower disk-based recovery places it slightly below LowDiff.
+For example, when the MTBF is set to 0.3 hours, LowDiff+ (S) leads with an effective training time ratio of 94.0%, followed closely by LowDiff at 92%.
+LowDiff+ (P) achieves 86.8%, which is still significantly higher than Gemini (81%) and Checkfreq (75.9%).
+These results highlight the robustness of our approaches, with LowDiff+ (S) offering the best performance for common software failures.
+Figure 18: (Exp. 9) Training under frequent failures.
 Experiment 10 (Scaling to more GPUs).
 We evaluate the training performance by varying the number of GPUs used in the training process, using the same effective training time ratio metric as in Experiment 7.
-As the number of GPUs increases, the probability of failures in the cluster also rises [19].
+As the number of GPUs increases, the probability of failures in the cluster also rises [gupta2024just].
 We conduct experiments with 8, 16, 32, and 64 GPUs, respectively, and measure the effective training time ratio of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
@@
 This highlights the strong potential of LowDiff and LowDiff+ in large-scale training.
-Figure 17: (Exp. 9) Training under frequent failures.
 Figure 19: (Exp. 10) Training with different number of GPUs.
-VIII Conclusion and Future Work
+IX Conclusion
 LowDiff is an efficient frequent checkpointing framework that addresses the challenges of the high costs of differential checkpointing by reusing the compressed gradients generated during training.

### v3

@@
 Recently, differential checkpointing has been proposed to reduce costs, but it is limited to recommendation systems, so its application to general distributed training systems remains unexplored.
-We proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost.
+We propose LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost.
 Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently.
 It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance.
-In non-compression scenario, We further proposes LowDiff+ with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression.
-Experiments on various workloads show that LowDiff can achieve checkpointing frequency up to per iteration with less than 3.1% runtime overhead.
+To enhance LowDiff under non-compression scenarios, we further propose LowDiff+, which incorporates a layer-wise-reuse snapshotting strategy, along with an incremental-merging persistence strategy.
+Experiments on various workloads show that LowDiff and LowDiff+ can reduce the training time by up to 89.2% and 81.2% with checkpointing frequency up to per iteration.
 I Introduction
@@
 Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [zhang2020empirical].
-For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [jeon2019analysis]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [OPT].
+For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [jeon2019analysis]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during a 90-day training period, averaging nearly two failures per day [OPT].
 To handle failures, checkpointing, which regularly writes the model state to storage (e.g., disks) for failure recovery, can confine the potential loss of the training progress to the interval between consecutive checkpoints, only requiring the recovery of the remaining progress.
@@
 Based on the findings, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials).
-Further, to enhance LowDiff, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
-In scenarios without gradient compression, we further investigate how to efficiently reuse gradients for frequent DC construction and propose LowDiff+.
-In this case, directly compressing differentials would incur high computation costs, so we reuse and snapshot gradients as they are generated layer by layer during the backward pass, overlapping their offloading with computation, thus enabling a fine-grained reusing and snapshotting approach.
-Meanwhile, the model state is reconstructed in CPU memory by the reused gradients and regularly persisted to storage, both in parallel with ongoing GPU training, thus enabling high-frequency in-memory checkpointing and asynchronous persistence.
+Further, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
+We further investigate how to reuse gradients for frequent checkpointing in scenarios without gradient compression.
+We first identify the issues of snapshotting and persistence when dealing with uncompressed gradients.
+To address these, we offer two key insights: (1) gradients can be reused layer-by-layer, and (2) differential and full checkpoints can be merged together.
+Based on the two insights, we proposed LowDiff+, which employs a layer-wise-reuse snapshotting strategy, along with an incremental-merging persistence strategy, thus enabling high-frequency checkpointing without gradient compression.
 Our contributions include:
@@
   • We perform an analysis of LowDiff, showing that gradient reuse eliminates the data dependency and enables parallel execution between checkpointing and training.
-Through measurement, we further demonstrate that checkpointing can be overlapped with the training iteration, allowing checkpointing to execute concurrently with training (§IV).
+Through measurement, we further demonstrate that checkpointing time can be overlapped with the training iteration, allowing checkpointing to execute concurrently with training (§IV).
   • We design LowDiff, which realizes the reuse of compressed gradients via a queue structure and facilitates zero-copy transmission between the training and checkpointing processes.
@@
   • We further propose LowDiff+ for efficient gradient reuse without compression.
-By exploiting the layer-by-layer generation of gradients during training, we design a layer-wise gradient reusing and snapshotting approach.
-We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§VI).
+By exploiting the layer-by-layer generation of gradients during training, LowDiff+ reuses and snapshots gradient in layer-wise granularity.
+LowDiff+ also merge the full model state and differentials in CPU memory before persistence to eliminates the overhead of writing separate differentials to storage (§VI).
   • We implement LowDiff and LowDiff+ atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff.
-Experiments on various models show that LowDiff and LowDiff+ can achieve checkpointing as frequently as once per iteration with less than minimal runtime overhead.
+Experiments on various models show that LowDiff and LowDiff+ can achieve checkpointing as frequently as once per iteration with minimal runtime overhead.
 Compared to the state-of-the-art methods, LowDiff and LowDiff+ can shorten training time by up to $89.2\%$ and $81.7\%$ under high checkpointing frequencies (§VII & §VIII).
@@
 $C^{D}_{t}=M_{t+1}-M_{t}.$ (5)
-Then the model state $M_{t}$ at the $t$-th iteration can be reconstructed by merging the full and all differential checkpoints as
+Then the model state $M_{t}$ at the $t$-th iteration can be reconstructed by merging the full and differential checkpoints as
 $M_{t}=C^{F}+C^{D}_{1}+C^{D}_{2}+\dots+C^{D}_{t-1}.$ (6)
@@
 IV Analysis
-By incorporating the main idea of gradient reuse (§III), LowDiff aims to enable the DC process to run concurrently with model training, thereby minimizing potential stalls.
+By incorporating the main idea of gradient reuse (§III), LowDiff aims to enable the checkpointing process to run concurrently with model training, thereby minimizing potential stalls.
 To confirm the feasibility of such parallelism, we analyze our proposed strategy from two key aspects:
-  • Parallelism property: it ensures no data dependencies block concurrency with training iterations.
-  • Execution time overlap: it reveals whether the checkpointing cost can be effectively hidden within the training time.
+  • Parallel of Checkpointing: it ensures no data dependencies block concurrency with training iterations.
+  • Overlap of execution time: it reveals whether the checkpointing cost can be effectively hidden within the training time.
 We provide a detailed analysis of these two aspects as follows.
-Figure 3: LowDiff’s DC can run in parallel with forward pass(F), backward pass(B), and model update(U).
-IV-A Parallelism Analysis
+Figure 3: LowDiff can run in parallel with forward pass ($F$), backward pass ($B$), and model update ($U$).
+IV-A Parallel of Checkpointing
 Necessity of Parallelism Analysis.
-Checkpointing can be executed with the forward and backward passes which do not modify the model state [mohan2021checkfreq], but it cannot run in parallel with model update which directly writes to it [wang2024fastpersist].
+Checkpointing can be executed with the forward and backward passes, which do not modify the model state [mohan2021checkfreq], but it cannot run in parallel with model update, which directly writes to it [wang2024fastpersist].
 In other words, checkpointing is data-dependent on the model update [wang2024fastpersist].
 The reason is that it must read the model state after model updates (Read-after-Write [hennessy2011computer], RAW dependency), and the model update at the next iteration must write the model state after checkpointing completes its read (Write-after-Read [hennessy2011computer], WAR dependency).
-Consequently, the two dependency patterns incur training stalls, preventing from paralleling checkpointing with model update.
+Consequently, the two dependency patterns incur training stalls, preventing parallelism of checkpointing with model update.
+To visualize such data dependency, Figure 3(a) shows the execution timeline with checkpointing.
+We see that the first dependency on $M_{t}$ (shown by the arrows from $U$ to $C$) and the second dependency on $M_{t}$ (shown by the arrows from $C$ to $U$) create a stall ($S$) before model update $U$ of the $i+1$-th iteration.
 Prior studies [mohan2021checkfreq, wang2023gemini, wang2024fastpersist] design their checkpointing strategies based on these data dependencies to pipeline training and checkpointing.
-Consequently, it is necessary to carefully analyze: (1) the data dependency introduced by our reuse strategy and (2) the parallelism property of LowDiff’s DC, specified as follows.
-(1) LowDiff’s DC can eliminate the WAR dependency. We next analyze current DC’s WAR dependency with next iteration’s model update, backward, and forward pass.
-  • Model update: as described in §II-A, the model update (Equation (4)) only requires read-only access to the gradient.
+Consequently, it is necessary to carefully analyze the data dependency introduced by our reuse strategy and the parallelism of checkpointing in LowDiff, specified as follows.
+LowDiff can eliminate the WAR dependency. We next analyze the WAR dependency with the next iteration’s model update, backward, and forward pass.
+  • Model update.
+As described in §II-A, the model update (Equation (4)) only requires read-only access to the gradient.
 Because LowDiff checkpoints only the compressed gradient, leveraging gradient reuse, the original Write-after-Read (WAR) dependency is eliminated because the checkpointing process no longer needs to read the model state ($M_{t}$).
@@
 Since both the checkpointing and model update are read-only consumers of the gradient, they can operate in parallel without conflict.
-  • Backward pass: LowDiff’s gradient reuse strategy introduces no additional WAR dependency on backward pass, since each gradient ${G}_{t}$ is finalized within its iteration and never modified thereafter.
+  • Backward pass.
+LowDiff’s gradient reuse strategy introduces no additional WAR dependency on backward pass, since each gradient ${G}_{t}$ is finalized within its iteration and never modified thereafter.
 Specifically, once a gradient ${G}_{t}$ is generated in the current iteration ($t$), it is final and will not be modified by subsequent steps.
-The backward pass of the $t+1$-th iteration (Equation (2)) does not read or depend on ${G}_{t}$; it generates a completely new gradient ($G_{t+1}$) from scratch and won’t change the ${G}_{t}$.
-Therefore, the DC’s read of ${G}_{t}$ does not conflict with the $B$’s write of ${G}_{t+1}$ at $t+1$-th iteration, allowing LowDiff’s DC to run in parallel with the next iteration’s backward pass.
-  • Forward pass: The forward pass (Equation (1)) doesn’t involve gradient, so there is no WAR dependency between LowDiff’s DC and forward pass.
-(2) LowDiff can run in parallel with forward pass, backward pass, and model update.
-To visualize the parallelism benefit from data dependency, Figure 3 shows the execution timeline with LowDiff.
-We see that the DC task ($C^{D}_{t}$) at the $t$-th iteration is no longer coupled with the model update ($U$) or the following iteration (no dependency arrow from DC to training).
+The backward pass of the $(t+1)$-th iteration (Equation (2)) does not read or depend on ${G}_{t}$; it generates a completely new gradient $G_{t+1}$ from scratch and won’t change ${G}_{t}$.
+Therefore, the DC’s read of ${G}_{t}$ does not conflict with the backward pass’s write of ${G}_{t+1}$ at the $(t+1)$-th iteration, allowing LowDiff to run in parallel with the next iteration’s backward pass.
+  • Forward pass. The forward pass (Equation (1)) doesn’t involve the gradient, so there is no WAR dependency between LowDiff and the forward pass.
+LowDiff can run in parallel with forward pass, backward pass, and model update.
+To visualize the parallelism benefit from data dependency, Figure 3(b) shows the execution timeline with LowDiff.
+We see that the checkpointing ($C_{t}$) at the $t$-th iteration is no longer coupled with the model update ($U$) or the following iteration (no dependency arrow from checkpointing to training).
 Instead, the reusing operation on compressed gradient ($\widetilde{G}_{t}$) begins immediately after the backward pass ($B$), running in parallel with the model update and subsequent iteration’s forward pass ($F$), backward pass ($B$), and model update ($U$).
-This parallelism property effectively transforms LowDiff’s DC from a blocking operation to a parallel execution with training, thereby eliminating the synchronization stalls previously caused by the WAR dependency.
+This parallelism property effectively transforms LowDiff from a blocking operation to a parallel execution with training, thereby eliminating the synchronization stalls previously caused by the WAR dependency.
 Figure 4: Time of iteration, full checkpointing, and differential checkpointing.
@@
 Beyond its parallelism property, LowDiff can perform even better if the checkpointing process is fast enough to be hidden by the training iteration, thus minimizing DC’s cost.
-Specifically, the time required to checkpoint a compressed gradient (denoted as $C$) should be shorter than a single training iteration (denoted as $F+B+U$), similar to the Microsoft study [mohan2021checkfreq].
-Consequently, it is also necessary to analyze the potential of overlapping, so we first (1) take a measurement analysis of training time, full checkpointing, and DC time, as well as (2) discuss why the overlap can be practically feasible.
+In other words, the time required to checkpoint a compressed gradient (denoted as $C$) should be shorter than a single training iteration (denoted as $F+B+U$), similar to the Microsoft study [mohan2021checkfreq].
+Consequently, it is also necessary to analyze the potential of overlapping, so we first (1) conduct a measurement analysis of training time, full checkpointing, and DC time, as well as (2) discuss why the overlap can be practically feasible.
 (1) The checkpointing of compressed gradients can be overlapped with training.
-Following the measurement setup in §III-A, we measure the iteration time, full checkpointing time and differential checkpointing time on BERT-B, BERT-L, GPT2-S, and GPT2-L.
+Following the measurement setup in §III-A, we measure the iteration time, full checkpointing time, and differential checkpointing time on BERT-B, BERT-L, GPT2-S, and GPT2-L.
 As shown in Figure 4, the DC time is only 20.5%, 22.1%, 23.1%, and 24.6% of the corresponding training iteration time for these models.
@@
 This duration is significantly shorter than that of a full training iteration [wang2023gemini].
-Based on the above analysis, we confirm that the concurrent execution between LowDiff’s DC and training is theoretically feasible.
-While this analysis establishes the feasibility of frequent DC, the detailed design of LowDiff will be presented in §V.
+Based on the above analysis, we confirm that the overlap of concurrent executions of LowDiff’s DC and training is practically feasible.
 V Design
@@
 Training first generates the local gradient, compresses it in Compressor, and synchronizes it across workers via Communicator.
-The synchronized compressed gradient $\widetilde{G}_{t}$ is put into Reusing Queue via zero-copy technique for reuse in checkpointing and is also decompressed by Decompressor to update the model state.
+The synchronized compressed gradient $\widetilde{G}_{t}$ is enqueued into Reusing Queue via zero-copy technique for reuse in checkpointing and is also decompressed by Decompressor to update the model state.
 Training checkpoints the model state $M_{t}$ regularly as full checkpoints $C^{F}_{t}$.
@@
 Reusing with zero-copy optimization.
-To meet Requirement 2, we use the CUDA Inter-Process Communication (IPC) queue in PyTorch to implement a zero-copy mechanism.
+To meet Requirement 2, we use the CUDA Inter-Process Communication (CUDA IPC) queue in PyTorch to implement a zero-copy mechanism.
 This allows GPU memory to be shared between training and checkpointing without extra copying, significantly reducing communication overhead.
@@
 In PyTorch, this can be achieved with torch.multiprocessing.Queue(), where the inter-process communication creates an IPC handle when a tensor is enqueued and reopens it when dequeued.
-Algorithm 1 Reusing in LowDiff
+Algorithm 1 LowDiff
 1: Model state $M_{t}$ at iteration $t$, Gradient $G_{i,t}$ at worker $i$, Reusing queue $Q$, Full checkpoint $C^{F}_{t}$, Differential checkpoint $C^{D}_{n}$ at iteration $n$, Compressor CompComp, Decompressor $Comp^{-1}$.
@@
 3:   for all $t$ do
-4:    $G_{i,t}\leftarrow Backward(x_{t})$
+4:    $G_{i,t}\leftarrow Backward(x_{t})$ $\triangleright$ Backward Pass
 5:    $\widetilde{G}_{i,t}\leftarrow Comp(G_{i,t})$ $\triangleright$ Compress
@@
 10: function Checkpointing Process
-11:   while $\widetilde{G}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Reusing
-12:    $C^{D}_{t}\leftarrow Save(\widetilde{G}_{t})$ $\triangleright$ Diff CKPT
-13:   $C^{F}_{t}\leftarrow Save(M_{t})$ $\triangleright$ Full CKPT Regularly
-14: function Recovery Process
-15:   $M_{t}\leftarrow Load(C^{F}_{t})$ $\triangleright$ Load Full CKPT
-16:   for $j\leftarrow t$, …, $n$ do $\triangleright$ Restore to Latest
-17:    $\widetilde{G}_{j}\leftarrow Load(C^{D}_{j})$ $\triangleright$ Load Diff CKPT
-18:    $G_{j}\leftarrow Comp^{-1}(\widetilde{G}_{j})$ $\triangleright$ Decompress
-19:    $M_{j+1}\leftarrow M_{j}+Adam(G_{j})$ $\triangleright$ Diff Merge
-20:   return $M_{n}$
-Design details Algorithm 1 specifies the following details of reusing compressed gradients, including training, checkpointing, and recovery processes.
+11:   for all $t$ do
+12:    while $\widetilde{G}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Reusing
+13:      $C^{D}_{t}\leftarrow Save(\widetilde{G}_{t})$ $\triangleright$ Diff CKPT
+14:    $C^{F}_{t}\leftarrow Save(M_{t})$ $\triangleright$ Full CKPT Regularly
+15: function Recovery Process
+16:   $M_{t}\leftarrow Load(C^{F}_{t})$ $\triangleright$ Load Full CKPT
+17:   for $j\leftarrow t$, …, $n$ do $\triangleright$ Restore to Latest
+18:    $\widetilde{G}_{j}\leftarrow Load(C^{D}_{j})$ $\triangleright$ Load Diff CKPT
+19:    $G_{j}\leftarrow Comp^{-1}(\widetilde{G}_{j})$ $\triangleright$ Decompress
+20:    $M_{j+1}\leftarrow M_{j}+Adam(G_{j})$ $\triangleright$ Diff Merge
+21:   return $M_{n}$
+Design details. Algorithm 1 specifies the following details of reusing compressed gradients, including training, checkpointing, and recovery processes.
 Training process.
@@
 Following synchronization, the training process puts the compressed gradient $\widetilde{G}_{t}$ into Reusing Queue $Q$ (Line 6), which will be transmitted to the checkpointing process with zero-copy.
-Finally, the training process decompresses the gradient (Line 6), updates the model state based on the gradient (Line 8), and proceeds to the next iteration.
+Finally, the training process decompresses the gradient (Line 7), updates the model state based on the gradient (Line 8), and proceeds to the next iteration.
 Checkpointing process.
@@
 Subsequently, the checkpointing process determines whether to merge the compressed gradient into batched differential checkpoints (detailed in §V-B).
-After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12-14) to storage.
-The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 15).
+After that, it invokes torch.save to persist the compressed differential checkpoints $C^{D}_{t}$ (Lines 12) to storage.
+The model state is also regularly checkpointed as a full checkpoint $C^{F}_{t}$ at the $t$-iteration (Line 13).
 Recovery process.
-During recovery, the recovery process first loads the latest full checkpoint $C^{F}_{t}$ at the $t$-th iteration to restore the model state $M_{t}$ (Line 17).
-Then, the compressed gradient $\widetilde{G}_{j}$ is retrieved from the differential checkpoint $C^{D}_{j}$ at the $j$-th iteration (Line 20) and decompressed into gradient $G_{j}$ (Line 21).
-Subsequently, the model state is restored from $M_{j}$ to $M_{j+1}$ at the $(j)$-th iteration by the optimizer and gradient (Line 21).
-By repeating this process for $n$ iterations (Lines 18-23), where compressed gradients from differential checkpoints are successively loaded and applied, the model state is finally restored to $M_{n}$ (Line 24), thus recovering the training from failures.
+During recovery, the recovery process first loads the latest full checkpoint $C^{F}_{t}$ at the $t$-th iteration to restore the model state $M_{t}$ (Line 15).
+Then, the compressed gradient $\widetilde{G}_{j}$ is retrieved from the differential checkpoint $C^{D}_{j}$ at the $j$-th iteration (Line 17) and decompressed into gradient $G_{j}$ (Line 18).
+Subsequently, the model state is restored from $M_{j}$ to $M_{j+1}$ at the $(j)$-th iteration by the optimizer and gradient (Line 19).
+By repeating this process for $n$ iterations (Lines 16), where compressed gradients from differential checkpoints are successively loaded and applied, the model state is finally restored to $M_{n}$ (Line 20), thus recovering the training from failures.
 V-B Batched Gradient Writing Optimization
@@
 The reason is that a low BS still incurs many writes due to insufficient batching, while a high BS makes one failure cause many gradients within the batch to be lost.
-Therefore, we need to tune both FCF and BS to minimize LowDiff’s wasted time. For example, the wasted time in Table I achieves its minimum when $FCF=20$ and $BS=2$.
+Therefore, we need to tune both FCF and BS to minimize LowDiff’s wasted time. For example, the wasted time in Table I achieves its minimum wasted time when $FCF=20$ and $BS=2$.
 FCF BS 1 2 3 4 5 6
@@
   • Total steady state overhead for checkpointing across all GPUs = $N\times\frac{S}{W}\times f\times T$.
-The wasted time $T_{wasted}$ can be calculated by adding recovery overhead and steady state overhead (see §II-B):
-$T_{wasted}$ $=\frac{NT}{M}(\frac{b}{2}+(R_{F}+\frac{R_{D}}{2}(\frac{1}{f\times b}-1)))+\frac{NT\times S\times f}{W}.$ (8)
-To minimize $T_{wasted}$ on two variables $f$ and $b$, we take the first-order partial derivatives of $T_{wasted}$ with respect to $f$ and $b$.
-$\begin{cases}\frac{\partial T_{wasted}}{\partial f}=\frac{NST}{W}-\frac{NT\times R_{D}}{2f^{2}\times M\times b}=0.\\ \frac{\partial T_{wasted}}{\partial b}=\frac{1}{2}-\frac{R_{D}}{2b^{2}\times f}=0.\end{cases}$ (9)
+The wasted time $T_{w}$ can be calculated by adding recovery overhead and steady state overhead (see §II-B):
+$T_{w}$ $=\frac{NT}{M}(\frac{b}{2}+(R_{F}+\frac{R_{D}}{2}(\frac{1}{f\times b}-1)))+\frac{NT\times S\times f}{W}.$ (8)
+To minimize $T_{w}$ on two variables $f$ and $b$, we take the first-order partial derivatives of $T_{w}$ with respect to $f$ and $b$.
+$\begin{cases}\frac{\partial T_{w}}{\partial f}=\frac{NST}{W}-\frac{NT\times R_{D}}{2f^{2}\times M\times b}=0.\\ \frac{\partial T_{w}}{\partial b}=\frac{1}{2}-\frac{R_{D}}{2b^{2}\times f}=0.\end{cases}$ (9)
 Equation (4) can be reduced to obtain the optimal values of the full checkpointing frequency ($f^{*}$) and batching size ($b^{*}$) as follows.
@@
 VI Enhancement
-The current design of LowDiff relies on gradient compression to enable efficient gradient reuse.
-However, in some high-precision training scenarios, any potential accuracy degradation caused by compression is unacceptable.
-If the distributed training system does not inherently employ a gradient compression strategy, making it impossible to leverage compression-induced bandwidth advantages.
-So we also need to design a non-compression high-frequency DC strategy for above scenarios.
-We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
-Figure 7: Layer-wise gradient reuse.
-VI-A Layer-Wise Gradient Reusing and Snapshotting Strategy.
-Issue of snapshotting.
-A key issue in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of a stall.
-The snapshotting competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall.
-To accelerate such large-scale snapshotting, we aim to design a fine-grained strategy that better overlaps gradient transmission with training.
+While LowDiff relies on gradient compression to achieve efficient gradient reuse, such compression may not be feasible or desirable in all training scenarios.
+In high-precision training tasks, gradient compression can introduce accuracy degradation, which is often unacceptable.
+Moreover, some distributed training systems may not employ gradient compression at all, thereby eliminating the bandwidth benefits that compression provides.
+To address these limitations, it is necessary to develop a high-frequency differential checkpointing (DC) strategy that does not depend on gradient compression.
+In this section, we first outline two key insights regarding snapshotting and persistence mechanisms in the absence of compression.
+We then present LowDiff+, an enhanced version of LowDiff tailored for scenarios where gradient compression is unavailable or undesired.
+(a) Snapshotting uncompressed gradient causes bandwidth contention and additional iteration time
+(b) Persisting large volume batched gradients requires multiple iteration time
+Figure 7: Issues of snapshotting and persistence.
+VI-A Issues of Snapshotting and Persistence
+Snapshotting issue.
+A key challenge in the non-compression setting arises during the snapshot phase: transferring uncompressed raw gradients from GPU to CPU introduces significantly heavier data movement compared to their compressed counterparts.
+This increased data traffic intensifies resource contention for GPU and I/O bandwidth, directly competing with training computation and substantially raising the risk of training stalls.
+To quantify the impact of bandwidth contention, we measure the iteration time with and without snapshotting of uncompressed gradients, following the experimental setup in §III-A.
+As shown in Figure 7a, snapshotting uncompressed gradients increases the iteration time by 12.3%, 16.7%, 13.1%, and 17.4% for BERT-B, BERT-L, GPT-S, and GPT-L, respectively.
+In contrast, snapshotting compressed gradients incurs only a 2.4%–3.1% overhead.
+This substantial performance gap highlights the need for optimized overlap mechanisms to mitigate training stalls.
+To accelerate large-scale snapshotting without compression, we need to design a fine-grained, layer-wise gradient reuse strategy that better overlaps gradient transmission with ongoing training.
+Figure 8: (Insight 1) Layer-wise gradient reuse.
+Figure 9: (Insight 2) Batching strategies comparison.
+Persistence issue.
+Although the layer-wise gradient reuse strategy mitigates the snapshotting overhead, a critical issue remains in the persistence phase when gradients are not compressed.
+The batched persistence strategy introduced in §V-B accumulates small compressed updates before writing them to storage, which is effective when updates are sparse and compact.
+However, without compression, differential checkpoints are substantially larger, and simply applying the same batching mechanism still demands write throughput that can easily saturate available storage bandwidth.
+[…]
+Figure 10: Architecture of LowDiff+.
+VI-B Insights into Snapshotting and Persistence
 Insight 1: Gradient can be reused layer-by-layer.
-We recall that gradients are produced incrementally in the backward pass, following the reverse layer order during training, as shown in Figure 7.
+It is found that gradients are produced incrementally in the backward pass, following the reverse layer order during training [rumelhart1986learning].
 Instead of waiting until the entire backward pass finishes to synchronize all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [deepspeed], PyTorch DDP [li2020pytorch], and Horovod [sergeev2018horovod]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced.
-Inspired by this execution model, we design a fine-grained reuse strategy: once a layer’s gradient has been synchronized, we immediately reuse it to the reusing queue, snapshot it to host memory, and stream it into a differential checkpoint assembler.
-This granular approach enables parallel data movement, allowing concurrent snapshot writes to CPU DRAM to better utilize memory bandwidth, thereby mitigating the high transfer volume.
-By pipelining backpropagation, communication, gradient reuse, and snapshotting at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
-Figure 8: Batching strategies of LowDiff and LowDiff+.
-VI-B CPU-Based Asynchronous Persistence Strategy.
-Issue of persistence.
-While the layer-wise strategy addresses the snapshotting issue in the non-compression setting, another issue remains: persistence of both full and differential checkpoints.
-The batching strategy described in §V-B accumulates small compressed updates on the CPU side before writing them out.
-It also periodically persists the full model state, which works effectively when update sizes are small and sparse.
-In the non-compression setting, differential checkpoints are significantly larger, and naively batching them demands a sustained write throughput that exceeds the available storage bandwidth, making it essential to design a more efficient checkpoint merging strategy that better lowers the writing data volume.
-Insight 2: differential and full checkpoints can be batched together in CPU memory.
-We find that in the non-compression setting, the differential checkpoints (i.e.
-gradients) and full checkpoints do not need to be persisted separately.
-Instead, they can be directly fused with the full model state inside CPU memory, enabling the CPU to maintain an always-up-to-date replica of the GPU model.
-Once gradients are merged into the CPU-resident model state, persistence only needs to occur from the CPU model to storage.
-By fusing gradients into the CPU-resident model state instead of persisting them separately, redundant writes of differential checkpoints are avoided, which directly reduces the overall storage volume.
-The different batching strategies of LowDiff and LowDiff+ are shown in Figure 8.
-This strategy decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
+Consequently, once a layer’s gradient has been synchronized, we can immediately reuse it in the reusing queue and snapshot it to host memory.
+This insight motivates a layer-wise fine-grain reuse and snapshotting design.
+As shown in Figure 8, taking the 4-th layer as an example, once its local gradient $G^{4}_{i,t}$ is computed (Equation (2)) and synchronized (Equation (3)), the resulting global gradient $G_{t}^{4}$ is immediately enqueued for reuse and snapshotting.
+Crucially, this process forms a cascading pipeline: as the backward pass propagates to earlier layers (3, 2, and 1), their computation and synchronization overlap with the snapshotting of later layers.
+This fine-grained interleaving maximizes memory bandwidth utilization and effectively hides the checkpointing overhead within the training execution, thereby minimizing stalls.
+Insight 2: Differential and full checkpoints can be merged incrementally.
+We observe that in LowDiff, differential checkpoints (i.e., raw gradients) are currently persisted separately from full checkpoints.
+Instead, they can be incrementally merged into the CPU-resident full checkpoint following the model update formula (Equation (4)), allowing the CPU to maintain an up-to-date replica of the GPU model.
+This merging strategy significantly reduces the volume of data to be persisted, as only the full model state, rather than individual differentials, needs to be persisted.
+This insight motivates an incremental merging design.
+As illustrated in Figure 9, LowDiff persists both differential ($C^{B}_{t}$) and full ($C^{F}_{t}$) checkpoints into storage.
+In contrast, we can maintain a CPU-resident model replica ($M_{t}^{C}$).
+The model state in CPU ($M_{t}^{C}$) is initialized only once from the full model state in GPU ($M_{t}^{G}$) at the start of training.
+During subsequent iterations, incoming differential checkpoints ($C^{D}_{t}$) are immediately applied to update $M_{t}^{C}$ in memory, keeping it synchronized with the GPU model.
+Consequently, only the full model state in CPU ($C^{F}_{t}$) requires persistence.
+By merging the differentials into the CPU-resident state incrementally, this insight eliminates the overhead of writing separate differential checkpoints to storage.
 VI-C Design of LowDiff+
-Design details.
-Based on Insight 1 and 2, we design LowDiff+ under non-compression scenarios.
-The architecture of LowDiff+ is shown in Figure 9, and its detailed algorithm is presented in Algorithm 2.
-In the training process, the model first performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer as shown in Figure 7 (Line 5).
-The layer-wise gradient communication is then handled by a communication thread pool $P_{g}$, where the gradient synchronization is assigned to multiple communication threads (Line 15).
-Once the communication in a thread finishes, the corresponding gradients are forwarded via the reusing queue to the CPU-side checkpointing process (Line 16), while the training process waits for the full gradient $G_{t}$ (i.e., wait for handle set $H_{g}$) (Line 6) to update the GPU model $M^{G}_{t}$ (Line 7).
-In the checkpointing process, gradients $G^{n}_{t}$ streamed from the GPU are reused layer by layer (Line 9).
-The offloading thread pool $P_{s}$ handles multiple layers in parallel (Line 10), snapshotting the layer-wise gradients into CPU memory as they become available (Line 19).
-This parallel snapshotting allows different layer gradients to be written concurrently to CPU memory, thereby better utilizing the available DRAM bandwidth.
-We initialize a CPU-resident replica $M^{C}_{t}$ of the model parameters, aligned with the GPU state.
-After snapshotting (Line 11), reused gradients are applied to $M^{C}_{t}$ to keep it synchronized with the GPU model $M^{G}_{t}$ (Line 12).
-While gradients are snapshotted layer-by-layer, the CPU-side update (Line 12) is performed once the full gradient is aggregated.
-This is because the Adam optimizer requires the complete gradient to correctly update its internal states (e.g., momentum and variance).
-The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [wang2023gemini].
-Since such updates do not involve forward or backward computations, their execution on the CPU introduces negligible overhead.
-Persistence is performed asynchronously by checkpointing the CPU-maintained model replica (Line 13), rather than the raw gradients.
-The frequency of persistence is decided similarly to Checkfreq [mohan2021checkfreq], which ensures maximal overlapping between the persist and update of the model in CPU.
-This design both avoids redundant storage of differential checkpoints and completely decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
-Figure 9: Architecture of LowDiff+.
-Recovery of LowDiff+.
-When failure occurs, prior work categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures.
-For hardware failures (e.g., GPU device crashes), the distributed training system must replace the failed machine with a new one.
-In this case, the in-memory training state is lost, so the recovery process must reload model states from persistent storage checkpoints, which can be time-consuming and increase wasted time.
-In contrast, software failures (e.g., software bugs) only cause the training process to terminate.
-This design assumes the failure is confined to the training processes (e.g., a CUDA error or NaN loss) and does not corrupt the separate checkpointing process’s memory, thus leaving the CPU memory state of the checkpointing process intact.
-In such cases, recovery is achieved by reinitializing the training process and restoring the preserved CPU-resident model state to the GPUs, leveraging in-memory checkpoints.
-This enables fast recovery under software failures, avoiding expensive checkpoint reloads and serving as a key optimization of LowDiff+.
-Algorithm 2 Reusing in LowDiff+
-1: GPU model state after the $t$-th iteration $M^{G}_{t}$, gradient of the $n$-th layer on node $i$ $G^{n}_{i,t}$, gradient communication and offload thread pools $P_{g}$ and $P_{s}$, reusing queue $Q$, corresponding handle sets $H_{g}$ and $H_{s}$.
+Algorithm 2 LowDiff+
+1: GPU model state $M^{G}_{t}$ at the $t$-th iteration, gradient $G^{n}_{i,t}$ of the $n$-th layer on node $i$ , reusing queue $Q$, CPU model state $M^{C}_{t}$.
 2: function Training Process
 3:   for all $t$ do
-4:    for all $n$ do
-5:      $G^{n}_{i,t}\leftarrow Backward(x^{n}_{t})$
-6:      $H_{g}.append(P_{g}.execute(\textsc{Sync},G^{n}_{i,t}))$ 7:                                         $\triangleright$ Sync thread
-8:    $G_{t}\leftarrow H_{g}.wait()$ $\triangleright$ Wait for sync
+4:    for all $n$ do $\triangleright$ Layer-by-Layer
+5:      $G^{n}_{i,t}\leftarrow Backward(x^{n}_{t})$ $\triangleright$ Backward Pass
+6:      $G^{n}_{t}\leftarrow Sync(G^{n}_{i,t})$ $\triangleright$ Synchronization
+7:      ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(G^{n}_{t})$ $\triangleright$ Gradient Reuse
+8:    $G_{t}\leftarrow\sum G^{n}_{t}$ $\triangleright$ Wait for Sync
 9:    $M^{G}_{t+1}\leftarrow M^{G}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in GPU
 10: function Checkpointing Process
-11:   while $G^{n}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Zero-copy reuse
-12:    $H_{s}.append(P_{s}.execute(\textsc{Snapshot},G^{n}_{t}))$ 13:                                       $\triangleright$ Snapshot thread
-14:   $G_{t}\leftarrow H_{s}.wait()$ $\triangleright$ Wait for offload
-15:   $M^{C}_{t+1}\leftarrow M^{C}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in CPU
-16:   $C^{F}_{t}\leftarrow Save(M^{C}_{t})$ $\triangleright$ Persistence fully
-17: function Sync Thread
-18:   $G^{n}_{t}\leftarrow Sync(G^{n}_{i,t})$ $\triangleright$ Synchronization
-19:   ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(G^{n}_{t})$ $\triangleright$ Zero-copy reuse
-20:   return $G^{n}_{t}$
-21: function Snapshot Thread
-22:   $G^{n}_{t}\leftarrow Offload\_to\_cpu(G^{n}_{t})$ $\triangleright$ Snapshot
-23:   return $G^{n}_{t}$
+11:   for all $t$ do
+12:    while $G^{n}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Gradient Reuse
+13:      $C^{n}_{t}\leftarrow Snapshot(G^{n}_{t})$ $\triangleright$ Snapshot
+14:    $G_{t}\leftarrow\sum C^{n}_{t}$ $\triangleright$ Wait for Snapshot
+15:    $M^{C}_{t+1}\leftarrow M^{C}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Merge in CPU
+16:    $C^{F}_{t}\leftarrow Persist(M^{C}_{t})$ $\triangleright$ Persist Regularly
+Based on Insights 1 and 2, we design LowDiff+ under non-compression scenarios. The architecture of LowDiff+ is shown in Figure 10, and its detailed algorithm is presented in Algorithm 2.
+Training Process.
+The model performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer (Line 4).
+Immediately following computation, each layer’s gradient undergoes synchronization (Line 5).
+Crucially, instead of blocking, the synchronized gradient $G^{n}_{t}$ is immediately enqueued into the reusing queue $Q$ (Line 6).
+This non-blocking operation allows the GPU to proceed with the backward pass of the next layer while the checkpointing process handles the data asynchronously.
+Finally, the training process aggregates the full gradient $G_{t}$ (Line 7) to update the GPU model state $M^{G}_{t+1}$ (Line 8).
+Checkpointing Process.
+The checkpointing process is structured into two stages: layer-wise-reuse snapshotting (based on Insight 1) and incremental-merging persistence (based on Insight 2).
+For layer-wise snapshotting, the checkpointing process continuously fetches gradients $G^{n}_{t}$ from the reuse queue $Q$ (Line 11) and performs an immediate snapshot to host memory (Line 12).
+For incremental-merging persistence, we maintain a CPU-resident replica $M^{C}_{t}$.
+After all layers’ gradients are snapshotted, the full gradient is aggregated in the CPU memory (Line 13).
+Once aggregated, the reused gradients are applied to $M^{C}_{t}$ to update its model state (Line 14).
+The updated model state serves as the in-memory checkpoint.
+Since this update executes asynchronously alongside the next training iteration in the GPU, its latency is effectively hidden.
+Finally, persistence is performed asynchronously by writing the updated CPU model replica to storage (Line 15).
 VII Implementation
@@
 Reusing queue module.
-Since Python’s GIL prevents true multi-threading, LowDiff uses torch.multiprocessing with the spawn method for process isolation.
-The compressed gradient Reusing Queue is implemented via torch.multiprocessing.Queue(), which (1) preserves checkpoint order and (2) leverages CUDA IPC for zero-copy GPU memory sharing.
+To bypass Python’s GIL, LowDiff uses torch.multiprocessing with the spawn method for process isolation.
+The Reusing Queue of compressed gradient is implemented via torch.multiprocessing.Queue(), which preserves checkpoint order and leverages CUDA IPC for zero-copy GPU memory sharing.
 Batched gradient writing module.
-Differential checkpoints are offloaded from GPU to CPU memory with offload_to_cpu() (non-blocking) and temporarily cached in a CPU buffer.
+Differential checkpoints are offloaded from GPU to CPU memory with Snapshot() (non-blocking) and temporarily cached in a CPU buffer.
 GPU memory handles are released after offloading.
@@
 Starting from a default setup, it adapts to runtime metrics (failure rate, bandwidth) using stepwise adjustments to balance steady-state overhead and recovery cost, achieving efficient fault-tolerant checkpointing.
-Parallel recovery module.Traditional recovery methods for differential checkpointing [eisenman2022check] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (6)), whose the number of merge operations is $n$ with given $n$ differential checkpoints.
+Parallel recovery module.Traditional recovery methods for differential checkpointing [eisenman2022check] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (6)), whose number of merge operations is $n$ with given $n$ differential checkpoints.
 Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints.
 Specifically, as shown in figure 11, it merges the earliest differential checkpoint $C_{1}^{D}$ with the full checkpoint $C_{t}^{F}$ while concurrently merging the subsequent differential checkpoints in pairs (i.e., $C_{2}^{D}$ and $C_{3}^{D}$, $C_{4}^{D}$ and $C_{5}^{D}$).
-This parallelization reduces the number of merge operations between differential checkpoint and full checkpoint from $n$ to $\log n$.
-In Figure 10, the number of merges between differential checkpoint and full checkpoint is 2 (the serial process has $n=5$ merges).
+This parallelization reduces the number of merge operations between the differential checkpoint and the full checkpoint from $n$ to $\log n$.
+In Figure 11, the number of merges between the differential checkpoint and the full checkpoint is 2 (the serial process has $n=5$ merges).
 Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
@@
 VII-B Implementation of LowDiff+
-LowDiff+ is also implemented atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed].
-Since the checkpointing process is spawned from the training process, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning.
-For gradient communication, we leverage Python’s concurrent.futures library and employ a ThreadPoolExecutor to assign each layer-wise gradient to a communication thread, ensuring parallel transmission.
-The snapshotting thread pool is implemented in the same manner, handling the writing of gradients into CPU memory.
-To maintain the CPU-side model replica, we configure PyTorch intra-op parallelism via torch.set_num_threads, setting the number of threads to match the available CPU cores in order to maximize parallel efficiency.
+LowDiff+ is also implemented atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed]. The implementation of the three main modules of LowDiff+ is as follows:
+Snapshotting module.
+To realize the layer-wise-reuse snapshotting execution, we leverage Python’s concurrent.futures library to manage concurrency through ThreadPoolExecutor.
+In the training process, we register backward hook for gradient synchronization, returning handles that allow the training loop to proceed non-blocking.
+On the checkpointing process, a separate thread pool concurrently retrieves gradients from the reusing queue and executes parallel snapshotting to the CPU DRAM.
+To prevent resource contention, we employ a priority-based policy that assigns a higher priority to gradient synchronization tasks over snapshotting tasks.
+This hierarchy ensures that in the event of bandwidth contention, the training task is always prioritized [zhang2025gockpt].
+Persistence module.
+To maintain the CPU-resident model replica, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning.
+we configure PyTorch intra-op parallelism via torch.set_num_threads.
 Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation.
-Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [mohan2021checkfreq], thereby maximizing checkpointing throughput without interfering with GPU training.
+Finally, the persistence to storage is performed asynchronously, following a design similar to Checkfreq [mohan2021checkfreq], which decouples disk I/O from the training critical path.
+Recovery module.
+When failure occurs, prior work [gupta2024just, wang2023gemini] categorizes failures into two types: hardware failures and software failures.
+For hardware failures (e.g., GPU crashes) where the training state is lost, the distributed training system must replace nodes and reload checkpoints from persistent storage, incurring significant downtime.
+Conversely, software failures (e.g., NaN loss) only cause the training process to terminate but leave the host operating system intact.
+Our design isolates the checkpointing process from the training process.
+Consequently, a software failure in the training process does not corrupt the separate memory space of the checkpointing process.
+Recovery is achieved by reinitializing the training process and directly restoring the preserved, up-to-date CPU-resident model state to the GPUs.
+The recovery from software failures bypasses expensive disk reads, enabling rapid recovery for the most common failure modes.
 VIII Evaluation
-In this section, we evaluate the efficiency of LowDiff to answer the following questions:
-  • How fast do LowDiff train under high checkpointing frequency compared to other methods? (Exp. 1)
-  • How fast do LowDiff+ train under non-compression scenarios? (Exp. 2)
-  • Can LowDiff and LowDiff+ achieve a shorter wasted time? (Exp. 3)
-  • Does LowDiff and LowDiff+ enable higher checkpointing frequency under bounded training speed? (Exp. 4)
-  • Does LowDiff and LowDiff+ accelerate the recovery process? (Exp. 5)
-  • Can LowDiff’s batched writes and offloaded batching reduce the checkpointing time and GPU memory cost? (Exp. 6)
-  • Can LowDiff reduce the storage overhead? (Exp. 7)
-  • Does LowDiff achieve frequent checkpointing under different compression ratios? (Exp. 8)
-  • Can LowDiff and LowDiff+ maintain high training efficiency under scenarios of more frequent failures and more GPUs? (Exp. 9-10)
 GPU Type GPU Mem CPU Type
@@
 TABLE II: Experimental setup
+In this section, we evaluate the efficiency of LowDiff to answer the following questions:
+  • How fast do LowDiff and LowDiff+ train under high checkpointing frequency with gradient compression? (Exp. 1)
+  • How fast do LowDiff and LowDiff+ train without gradient compression? (Exp. 2)
+  • Can LowDiff and LowDiff+ achieve a shorter wasted time? (Exp. 3)
+  • Does LowDiff and LowDiff+ enable higher checkpointing frequency under bounded training speed? (Exp. 4)
+  • Does LowDiff and LowDiff+ accelerate the recovery process? (Exp. 5)
+  • Can LowDiff’s batched writes and offloaded batching reduce the checkpointing time and GPU memory cost? (Exp. 6)
+  • Can LowDiff reduce the storage overhead? (Exp. 7)
+  • Does LowDiff achieve frequent checkpointing under different compression ratios? (Exp. 8)
+  • Can LowDiff and LowDiff+ maintain high training efficiency under scenarios of more frequent failures and more GPUs? (Exp. 9-10)
 VIII-A Experimental Setup
@@
 We also implement the LowDiff system on the VGG-16 model with pipeline parallelism provided in the DeepSpeedExamples [deepspeedexample].
-Baselines.
+Comparative Methods.
 We evaluate two state-of-the-art checkpointing strategies: CheckFreq [mohan2021checkfreq] and Gemini [wang2023gemini], as described in §II-B.
@@
 Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [eisenman2022check] does in recommendation systems.
+For LowDiff+, we denote the snapshotting phase as LowDiff+(S) and the persistence phase as LowDiff+(P).
 VIII-B Performance Evaluation
-Figure 11: (Exp. 1) Training time of different models with different checkpointing strategies.
-Experiment 1 (Training time of LowDiff).
-We evaluate the training time of LowDiff under different checkpointing strategies on the previously introduced DNN models, with the checkpointing frequency set to once per iteration.
+Figure 12: (Exp. 1) Training time with gradient compression.
+Experiment 1 (Training time with gradient compression).
+We evaluate the training time with gradient compression of LowDiff and LowDiff+ under different checkpointing strategies on the previously introduced DNN models, with the checkpointing frequency set to once per iteration.
 The experiment is conducted in the presence of gradient compression.
@@
 The W/O CKPT setting represents training without checkpointing, which serves as the upper bound of achievable training speed.
-Figure 11 shows that the training time of LowDiff on all seven training tasks is close to that of W/O CKPT and significantly smaller than those of other solutions.
-Compared to W/O CKPT, LowDiff’s training time increases by only 2.4%-3.1%, while the other checkpointing methods’ training time increases by 8.1%-891%.
-Specifically, we elaborate on LowDiff’s benefits over the other methods and analyze the reasons behind these benefits:
+Figure 12 shows that the training time of LowDiff and LowDiff+ on all seven training tasks is close to that of W/O CKPT and significantly shorter than that of other solutions.
+Compared to W/O CKPT, LowDiff and LowDiff+’s training time increases by only 2.4%-3.1% and 0.7%-2.9%, respectively, while the other checkpointing methods’ training time increases by 8.1%-891%.
+Specifically, we elaborate on LowDiff and LowDiff+’s benefits over the other methods and analyze the reasons behind these benefits:
   • Compared to CheckFreq and Gemini in GPT2-S, LowDiff reduces the training time by 68.2% and 46.1%, respectively.
@@
 Our future work will explore the efficiency of such a combination.
-Figure 12: (Exp. 2) Training time without compression.
-Experiment 2 (Training time of LowDiff+).
-To isolate the impact of gradient compression, we further evaluate the training time of LowDiff+ in a setting without gradient compression.
-All other experimental configurations, including the checkpointing frequency (per iteration), the selected DNN models, the number of training iterations (1,000), and the hardware environment (A100 servers), remain identical to Exp.
+  • For LowDiff+, its training time across all models is even shorter than LowDiff.
+Specifically, LowDiff+ reduces the training time by 0.8%, 1.5%, 1.7%, 1.6%, 1.1%, and 3.3% in ResNet-101, VGG-19, BERT-B, BERT-L, GPT2-S, and GPT2-L, respectively.
+While the compressed gradients are already small enough to be handled efficiently by LowDiff’s DC, limiting the gains in snapshotting and persistence of LowDiff+, the advantages of LowDiff+ become more pronounced in non-compression scenarios, as further discussed in Exp.
+2.
+Figure 13: (Exp. 2) Training time without gradient compression.
+Experiment 2 (Training time without gradient compression.) We further evaluate the training time of LowDiff and LowDiff+ in a scenario where gradient compression is disabled, while keeping other settings the same as Exp.
 1.
-Figure 12 shows that the training time of LowDiff+ across all tasks exhibits a slight increase compared to the W/O CKPT baseline, yet it remains significantly lower than that of other checkpointing solutions.
-Specifically, LowDiff+’s training time increases by 7.2%–9.1% relative to W/O CKPT.
-As expected, this overhead is higher than the 2.4%-3.1% observed for LowDiff in Exp.
-1.
-This increase represents the necessary and expected performance trade-off for handling uncompressed, high-precision gradients, which was the primary motivation for the LowDiff+ design.
-The overhead is mainly due to frequent and large-volume gradient transfers between GPU and CPU memory, occupying PCIe bandwidth.
-Despite this minor overhead, in the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods.
-For the GPT2-L training task, it reduces the training time by 51.8% compared to Gemini and by 81.7% compared to CheckFreq, demonstrating its superior efficiency in settings where frequent non-compression checkpoints are required.
+As shown in Figure 13, LowDiff+ consistently achieves the lowest training time among all evaluated checkpointing methods.
+Specifically, LowDiff+ consistently achieves the lowest training time among all checkpointing methods.
+For the GPT2-L training task, it reduces the training time by 42.4% compared to Gemini and by 81.2% compared to CheckFreq, demonstrating its efficiency in settings where frequent non-compression checkpoints are required.
+Furthermore, the results highlight the effectiveness of the proposed enhancements in LowDiff+.
+In this non-compression setting, the original LowDiff incurs significant overhead, with training time increasing by 16.1%–28.1% compared to the W/O CKPT baseline across different models.
+In contrast, LowDiff+ limits this overhead to an average of only 5.1%-7.2%.
+Although this is higher than the 1.8% overhead observed in Exp.
+1 (due to the increased bandwidth pressure from uncompressed gradients), the substantial improvement over LowDiff confirms that our layer-wise-resue snapshotting and incremental-merging persistence successfully mitigate the snapshotting and persistence issues in §VI-A.
+Figure 14: (Exp. 3) Wasted time under different MTBFs.
 Experiment 3 (Wasted time).
@@
 Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §V-C.
-For LowDiff+, we evaluate its wasted time under both software and hardware failures, where the corresponding checkpoints can be accessed throught snapshot and persist ( LowDiff+ (S) and LowDiff+ (P), respectively).
+For LowDiff+, we evaluate its wasted time under both software and hardware failures, where the corresponding checkpoints can be accessed through LowDiff+(S) and LowDiff+(P).
 Figure 14 shows that the wasted time gap between LowDiff and other checkpointing methods keeps increasing as failures become more frequent, while LowDiff maintains the lowest wasted time.
@@
 The results confirm the efficiency of LowDiff in the event of failures.
-For LowDiff+, the wasted time due to software failures is 3.7%–5.1% lower than that of LowDiff.
-This improvement comes from LowDiff+ ’s in-memory checkpointing, which enables fast recovery using the full model state stored in CPU memory.
-The detailed recovery times of LowDiff+ (S) are presented in Exp.
-5.
-In contrast, the wasted time from hardware failures is slightly higher than that of LowDiff, due to LowDiff+ ’s lower checkpointing frequency, but still lower than that of CheckFreq and Gemini.
-Details of the checkpointing frequency for LowDiff+ (S) and LowDiff+ (H) are presented in Exp.
-6.
-Figure 13: (Exp. 3) Wasted time under different MTBFs.
+[…]
 Experiment 4 (Checkpointing frequency).
@@
 Figure 15 shows that compared to other methods, LowDiff can achieve the checkpointing frequency every iteration (i.e., the checkpointing frequency equals one) for different widely-used models under the bounded training speed.
-LowDiff+ (S) also performs per-iteration in-memory checkpointing, enabled by its efficient layer-wise gradient reuse and snapshotting approach.
+LowDiff+(S) also performs per-iteration in-memory checkpointing, enabled by its efficient layer-wise-reuse snapshotting approach.
 LowDiff+ (P) achieves per-iteration persistence frequency in smaller models such as ResNet-101, but increases to 3 iterations for larger models like GPT2-L due to the limited bandwidth of SSDs.
-Nevertheless, its checkpointing intervals remain the shortest compared to Naïve DC, CheckFreq, and Gemini.
+Nevertheless, its checkpointing intervals remain shorter than those of Naïve DC and CheckFreq, and are nearly the same as those of Gemini.
 Intervals of Naïve DC grow significantly with model size, increasing from every 2 iterations to every 8.
@@
 Figure 15: (Exp. 4) Maximum checkpointing frequency.
-Figure 15: (Exp. 5) Recovery time of different methods.
 Experiment 5 (Recovery time).
-We evaluate the recovery time of different recovery strategies under different full checkpointing frequency in GPT2-S.
-We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [paszke2019pytorch]), Naïve DC, LowDiff with parallel recovery (see §VII), and LowDiff+ from software failures (LowDiff+ (S) for short).
+We evaluate the recovery time of different recovery strategies under different full checkpointing frequencies in GPT2-S.
+We compare traditional checkpointing without differential checkpoint as Baseline (i.e., Torch.save [paszke2019pytorch]), Naïve DC, LowDiff with parallel recovery (see §VII), and LowDiff+ from software failures (LowDiff+(S) for short).
 Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
 […]
+Figure 16: (Exp. 5) Recovery time of different methods.
 (a) Average checkpointing time under different batch sizes
@@
 For GPT2-L, LowDiff supports per-iteration checkpointing when $\rho\in[0.001,0.075]$, while at $\rho=0.1$, the checkpoint frequency increases to 2 iterations because the increasing gradient size makes the checkpointing take two iterations to be overlapped with training.
-The result shows LowDiff ensures frequent checkpointing (intervals ¡ 3 iterations) under the common compression ratio.
+The result shows LowDiff ensures frequent checkpointing (intervals <3 iterations) under the common compression ratio.
 VIII-C System Scalability
@@
 Using the effective training time ratio metric as defined in Gemini [wang2023gemini], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
-Note that LowDiff+ has two recovery modes: LowDiff+ (S) for software failures (recovering from CPU memory) and LowDiff+ (P) for hardware failures (recovering from persistent storage).
-Figure 18 demonstrates that LowDiff+ (S) consistently achieves the highest effective training time ratio across all scenarios, especially under environments with frequent failures.
-This is because its near-instantaneous recovery from in-memory checkpoints (as shown in Exp.
-5) almost completely eliminates recovery overhead, compensating for its slightly higher steady-state overhead compared to LowDiff.
+Figure 19 demonstrates that LowDiff+ consistently achieves the highest effective training time ratio across all scenarios, especially under environments with frequent failures.
+This is because its fast recovery from in-memory full model state (as shown in Exp.
+5) almost eliminates recovery overhead, compensating for its slightly higher steady-state overhead compared to LowDiff.
 LowDiff achieves the second-highest ratio.
-LowDiff+ (P) also consistently outperforms the other baselines, though its reliance on slower disk-based recovery places it slightly below LowDiff.
-For example, when the MTBF is set to 0.3 hours, LowDiff+ (S) leads with an effective training time ratio of 94.0%, followed closely by LowDiff at 92%.
-LowDiff+ (P) achieves 86.8%, which is still significantly higher than Gemini (81%) and Checkfreq (75.9%).
-These results highlight the robustness of our approaches, with LowDiff+ (S) offering the best performance for common software failures.
-Figure 18: (Exp. 9) Training under frequent failures.
+For example, when the MTBF is set to 0.3 hours, LowDiff+ leads with an effective training time ratio of 93.3%, followed closely by LowDiff at 92.1%.
+These results highlight the robustness of our approaches, with LowDiff+ offering the best performance for common software failures.
 Experiment 10 (Scaling to more GPUs).
@@
 We conduct experiments with 8, 16, 32, and 64 GPUs, respectively, and measure the effective training time ratio of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
-Figure 19 shows that, as the number of GPUs increases, the effective training time ratio decreases; however, LowDiff still maintains a 98% proportion of the effective training time ratio, and LowDiff+ achieves 96%, while the metric for the other method declines rapidly and can only reach around 90%.
+Figure 20 shows that, as the number of GPUs increases, the effective training time ratio decreases; however, LowDiff and LowDiff+ still maintain a nearly 98% proportion of the effective training time ratio, while the metric for the other method declines rapidly and can only reach around 90%.
 This highlights the strong potential of LowDiff and LowDiff+ in large-scale training.
+Figure 19: (Exp. 9) Training under frequent failures.
 Figure 20: (Exp. 10) Training with different number of GPUs.
@@
 To enhance the gradient reuse strategy in non-compression scenarios, we design LowDiff+.
-LowDiff+ uses a layer-wise gradient reusing and snapshotting strategy and a CPU-based synchronous persistence strategy, thus enabling checkpointing to decouple from training.
+LowDiff+ uses a layer-wise-reuse snapshotting strategy and an incremental-merging persistence strategy, thus enabling checkpointing to decouple from training.
 Our experiments on a variety of models and GPU generations demonstrate the efficiency of LowDiff and LowDiff+ in terms of the training time, wasted time, checkpointing time, and recovery time under the high-frequency checkpointing scenarios.